### v3

# Quantising Chiral Bosons On Riemann Surfaces

Abstract
Sen’s action in two dimensions governs a chiral boson coupled to a two-dimensional metric together with a second chiral boson that couples to a flat two-dimensional metric. This second scalar decouples from the physical degrees of freedom. The generalisation of this action to one in which the second chiral scalar couples to an arbitrary second metric is used to formulate the theory on an arbitrary two-dimensional manifold. We use this action with both metrics Riemannian (or complex) to formulate the path integral on any Riemann surface. We calculate the partition function in this way and check the result with that calculated using canonical quantisation, and then extend this to multiple chiral bosons. The partition function for chiral scalars taking values on a rational torus is a sum of terms, each of which is the product of two holomorphic functions, one a function of the modulus of the first metric and the other a function of the modulus of the second metric. In particular, for the case of chiral bosons moving on a torus defined by an even self-dual lattice, the partition function is a single product of two such holomorphic functions, not a sum of such terms. This is applied to the heterotic string to give a world-sheet action whose quantisation is modular invariant and free from anomalies. We discuss modular invariance for the moduli of both metrics and the extension to higher genus Riemann surfaces.
preprint: Imperial-TP-2025-CH-5
1 Introduction and Summary
In Sen:2015nph ; Sen:2019qit Sen gave an interesting action for $2k$-form gauge fields in $4k+2$ dimensions with self-dual field strength. The degrees of freedom of Sen’s theory consist of the physical $2k$-form gauge field $A$ with self-dual field strength $F$ that couples to the spacetime metric $g$ and the other physical fields, together with a second $2k$-form gauge field $C$ with field strength $G$ that doesn’t couple to the spacetime metric $g$ and the other physical fields. Instead, $C$ couples to a flat metric $\eta$ and $G$ is self-dual with respect to $\eta$. Thus the theory consists of the physical chiral gauge field $A$ plus a second chiral gauge field $C$ which decouples from the physical theory. The theory has a diffeomorphism-like symmetry that acts on $F$ but not $G$. The interaction with the space-time metric $g$ is governed by a tensor $M$ which was constructed perturbatively in Sen:2015nph ; Sen:2019qit . This theory has been further explored in Lambert:2019diy ; Andriolo:2020ykk ; Barbagallo:2022kbt ; Vanichchapongjaroen:2024tkj ; Aggarwal:2025fiq ; Hull:2025yww . Other approaches to chiral gauge fields are discussed in e.g. Siegel:1983es ; Floreanini:1987as ; Devecchi:1996cp ; Pasti:1996vs ; Belov:2006jd ; Mkrtchyan:2019opf ; Arvanitakis:2022bnr and references therein.
In Hull:2023dgp a generalisation of Sen’s action was constructed in which the non-physical gauge-field $C$ couples to an arbitrary second metric $\bar{g}$ instead of the flat metric $\eta$, and $G$ is self-dual with respect to $\bar{g}$. The theory then describes a physical sector consisting of the gauge field $A$, the spacetime metric $g$ and the other physical fields, together with a shadow sector consisting of the gauge-field $C$ and the second metric $\bar{g}$. The shadow sector decouples from the physical sector. This has a number of advantages over Sen’s formulation. First, it can be used on any spacetime manifold, not just those admitting a flat metric $\eta$. Second, it allows a geometric construction of the tensor $M$ in closed form (using results from Andriolo:2020ykk ). Indeed, finding the tensor $M$, which has a complicated dependence on both metrics $g$ and $\bar{g}$ was a non-trivial part of the the construction of Hull:2023dgp . Next, it has two gauge symmetries with vector parameters, one for which $g$ is the gauge field and one for which $\bar{g}$ is the gauge field. The diffeomorphism symmetry is a diagonal subgroup of these two symmetries. A related use of a bi-metric theory is proposed for string field theory in Hull:2025mtb .
Both Sen’s formulation and its bi-metric generalisation are quadratic in the dynamical fields and so are straightforward to quantise. The purpose of this paper is to explore the quantisation of a chiral scalar in 2 dimensions (the case $k=0$) using the action of Hull:2023dgp , with $A$ and $C$ chiral scalars. In particular, we calculate the partition function and show that it has the desired form, which is a holomorphic function of the modulus $\tau$ if the scalar is periodic. Similar calculations were performed before in the case of a torus where one can use the original Sen action with $\bar{g}=\eta$ Andriolo:2021gen ; Lambert:2023qgs .
The physical metric can be written in the form
$g=e^{2\omega}(dx+\tau dy)(dx+\tilde{\tau}dy).$ (1)
where $\tau$ and $\tilde{\tau}$ are independent real functions for Lorentzian signature and are conjugate complex functions for Euclidean signature. The action of Hull:2023dgp describes chiral bosons in Lorentzian signature. However, for the quantum theory we are interested in the path integral for Euclidean signature metrics. In particular, for the partition function we are interested in the case in which the 2-dimensional space is a 2-torus with metric $g$ (1) for which $\tau$ is a complex constant, the modulus, with complex conjugate $\tilde{\tau}$.
The shadow metric $\bar{g}$ can be written in the same form
$\bar{g}=e^{2\sigma}(dx+\rho dy)(dx+\tilde{\rho}dy).$ (2)
where as before $\rho$ and $\tilde{\rho}$ are independent real functions for Lorentzian signature and are conjugate complex functions for Euclidean signature. We will proceed by using the action of Hull:2023dgp with complex metrics $g$ and $\bar{g}$, so that the real Lorentzian and Euclidean metrics arise as special cases. Specifically, we will take the metrics $g$ and $\bar{g}$ to be given by (1) and (2), but will consider $\tau$ and $\tilde{\tau}$ and $\rho$ and $\tilde{\rho}$ to be independent complex variables. For a 2-torus, the symmetries can be used to set $\tau,\tilde{\tau}$ and $\rho,\tilde{\rho}$ to complex constants, while for higher genus they are functions on the Riemann surface that depend on the moduli. We will calculate in regions of the space of complex variables $\tau,\tilde{\tau},\rho,\tilde{\rho}$ for which the functional integral is well-defined and then analytically continue the results to the whole parameter space. Complex metrics have also recently been considered in the context of quantum calculations in e.g. Kontsevich:2021dmb ; Witten:2021nzp ; BenettiGenolini:2025jwe . The scalar fields remain real in our calculations.
In two dimensions, the M-tensor that incorporates all dependence on the metric $g$ has one independent component, which we will denote by $M_{--}$ in what follows. For a given metric $g$, we will choose $\bar{g}$ to be such that $M_{--}$ is finite and non-zero everywhere; this is helpful in calculating the path integrals. We will then analytically continue the result for general $g,\bar{g}$.
We find a strikingly elegant and simple result. For a theory of $8k$ chiral bosons associated with a torus $\mathbb{R}^{8k}/\Delta$ where $\Delta$ is an even self-dual $8k$-dimensional lattice, the partition function depends on just two of the four moduli and has the factorised form
$Z(\tau,\rho)=F(\tau)\tilde{F}(\rho).$ (3)
This has immediate application to the heterotic string, which involves 16 chiral bosons and an even self-dual lattice which is either the root lattice of $E_{8}\times E_{8}$ or the weight lattice of $Spin(32)/\mathbb{Z}_{2}$. Then the bi-metric chiral boson theory can be used for the heterotic string where further world-sheet fields coupling to $g$ cancel the modular anomaly coming from $F(\tau)$, so that the full theory is invariant under $\tau$ modular transformations. We will see how invariance under $\rho$ modular transformations could also be achieved by introducing further world-sheet fields coupling to $\bar{g}$.
More generally, for a chiral boson on a circle or for $n$ chiral bosons on a torus, the partition function again depends on just two of the four moduli, $\tau$ and $\rho$. In general it is of the form
$Z(\tau,\rho)=F_{o}(\tau)\tilde{F}_{o}(\rho)f(\tau,\rho).$ (4)
The contribution from the oscillators gives a factorised form $F_{o}(\tau)\tilde{F}_{o}(\rho)$ while the integration over the momenta and winding gives a function $f(\tau,\rho)$ which in general doesn’t factorise. For the special case in which the compactification defines a rational conformal field theory (e.g. for a boson on a circle of radius $R$ with $R^{2}$ rational) then $f(\tau,\rho)$ is a finite sum of terms each of which factorise, so that the partition function takes the form
$Z(\tau,\rho)\sim\sum_{a,\bar{b}}C_{a\bar{b}}\chi_{a}(\tau)\chi_{\bar{b}}(\rho),$ (5)
for some constants $C_{a\bar{b}}$ and some functions $\chi_{a}(\tau),\chi_{\bar{b}}(\rho)$ (which are Verma module characters).
This arises in an interesting way. A central role in our analysis is played by an effective action for a non-chiral scalar coupled to a complex metric
$h_{\mu\nu}dx^{\mu}dx^{\nu}$ $=e^{2\sigma^{\prime}}(dx+\rho dy)\left(dx+\tau dy\right),$ (6)
with moduli $\rho,\tau$. We quantise the scalar coupled to this complex metric by doing the partition function in regions of the $\rho,\tau$ moduli space in which the functional integral is well-defined and then analytically continuing.
Chiral bosons arise in a number of different situations. On 2-dimensional Minkowski space we can have chiral bosons which are non-compact (taking values on a line) or compact (taking values on a circle). For fields on a 2-dimensional cylinder, as arise in closed string theory, a chiral boson with non-trivial momentum can only be achieved for compact bosons, and moreover the radius $R$ of the target space circle must satisfy $R^{2}=1$ (in units in which the string tension is $1/2\pi$, i.e. $\alpha^{\prime}=1$). For multiple chiral bosons, they must take values in a rational torus. This will be discussed in detail later. However, the bi-metric chiral boson action is defined for any 2-dimensional space and any flat target space, either Euclidean space, or a torus which does not need to be rational, and the resulting partition function can have complicated dependence on both $\tau$ and $\rho$. However, for the case of a rational torus the theory simplifies and the partition function is a sum of a finite number of terms, each of which factorises into a function of $\tau$ times a function of $\rho$.
One of our main motivations is in the application to string theory where one would want to study chiral bosons on an arbitrary Riemann surface. The bi-metric action is well-suited for this and we study both the torus case and higher-genus Riemann surfaces here. We will discuss the applications to heterotic string theory correlators elsewhere.
The rest of this paper is organised as follows. In section 2 we discuss the action in detail and evaluate it in the special case in which the metrics $g$ and $\bar{g}$ are given by (1) and (2) respectively. In section 3 we turn to the quantum theory. We first discuss the canonical quantization, generalised to a complex metric, and obtain the partition function. We then compute in detail the path integral over a torus to find the same expression. In section 4 we discuss modular invariance and what the ranges of the parameters $\rho,\tilde{\rho},\tau,\tilde{\tau}$ must be to ensure convergence of the integrals in the path integral. In section 5 we extend our results to multiple chiral scalars and show how to obtain a partition function that factorises in to a simple product as in (4) and construct a consistent heterotic string theory by including matter which cancel anomalies in both diffeomorphism symmetries. In section 6 we evaluate the path integral on a generic Riemann surface. Finally, in section 7 we present our conclusions and a discussion of our results.
2 The Action
2.1 The Action For Chiral Gauge Fields
The Sen action Sen:2015nph ; Sen:2019qit is defined in $4k+2$ dimensions. It’s generalisation in the presence of a source $J$ and coupled to a general background metric is Hull:2023dgp
$S=-\int\Big($ $\frac{1}{2}dP\wedge\bar{\star}dP-2Q\wedge dP-(Q+J)\wedge{\cal M}(Q+J)$
$+2Q\wedge J-\frac{1}{2}J\wedge\bar{\star}J\Big).$ (7)
[…]
${\cal M}(Q)=\frac{1}{2}(1-\bar{\star}){\cal M}\left(\frac{1}{2}(1+\bar{\star})Q\right).$ (8)
It is also taken to be symmetric in the sense that
$Q_{1}\wedge{\cal M}(Q_{2})=Q_{2}\wedge{\cal M}(Q_{1}),$ (9)
[…]
The role of ${\cal M}(Q)$ is to ensure that
$Q+{\cal M}(Q)=\star\left(Q+{\cal M}(Q)\right),$ (10)
where $\star$ is the Hodge dual associated to a second, physical metric $g$ that is independent of $\bar{g}$. The equations of motion of the action (2.1) take the form
$0$ $=d\left(\bar{\star}dP+2Q\right)$
$0$ $=-dP+\bar{\star}dP-2{\cal M}(Q+J)+(1-\bar{\star})J.$ (11)
We define
$F$ $=Q+\tfrac{1}{2}(1+\bar{\star})J+{\cal M}(Q+J)$
$G$ $=Q+\frac{1}{2}(1+\bar{\star})dP,$ (12)
so that $F$ is self-dual with respect to $g$ while $G$ is self-dual with respect to $\bar{g}$:
$F=\star F,\qquad G=\bar{\star}G.$ (13)
The equations of motion can be rewritten as
$dG$ $=0$
$dF$ $=dJ.$ (14)
In this way, we find two dynamical form fields, one self-dual with respect to the one of the metrics and the other self-dual with respect to the other metric. However, only $F$ couples to the source and to the physical metric $g$. While the source $J$ is kept arbitrary here, in other examples it takes on an explicit form in terms of various matter fields, for example in terms of other form fields and fermions in type IIB supergravity.
2.2 The action For Chiral Scalars
In this paper we will focus on the simplest example of two-dimensions, where $P$ is a scalar and $Q$ a 1-form. It is helpful to use a zweibein frame where (see the appendix for our conventions and notation)
${\bar{g}}_{\mu\nu}$ $={\bar{e}}_{\mu}{}^{a}{\bar{e}}_{\nu}{}^{b}\eta_{ab}$
$={\bar{e}}_{\mu}{}^{+}{\bar{e}}_{\nu}{}^{-}+{\bar{e}}_{\mu}{}^{-}{\bar{e}}_{\nu}{}^{+}.$ (15)
In this case one sees that the self-duality condition on $Q$ simply becomes
$Q=Q_{+}\bar{e}^{+},$ (16)
so the 1-form $Q$ has only one non-vanishing component $Q_{+}(x)$. Furthermore, in such a basis we must have that
${\cal M}(Q_{+}\bar{e}^{+})=M_{--}Q_{+}\bar{e}^{-},$ (17)
so the symmetric tensor $M_{\mu\nu}$ has only one non-vanishing component, $M_{--}(x)$.
Rescaling the action (2.1) by the string tension we find
$S$ $=-\frac{1}{2\pi}\int{\rm det}(\bar{e})\Big(-\partial_{+}P\partial_{-}P-2Q_{+}\partial_{-}P-(Q_{+}+J_{+})^{2}M_{--}$
$\hskip 113.81102pt+2Q_{+}J_{-}+J_{+}J_{-}\Big)d^{2}x,$ (18)
where $\partial_{\pm}=\bar{e}_{\pm}{}^{\mu}\partial_{\mu}$. The field strengths with $J=0$ are
$G_{+}=\frac{1}{2}\partial_{+}P+Q_{+},\qquad F_{+}=Q_{+},\qquad F_{-}=M_{--}Q_{+},$ (19)
and $G_{-}=0$, $F=\ast F$.
As discussed in the introduction, for a given metric $g$ we will restrict our attention to choices of $\bar{g}$ so that $M_{--}(x)$ is well-defined and non-zero for all $x$. We then find that the action (15) can be rewritten as
$S$ $=-\frac{1}{2\pi}\int{\rm det}(\bar{e})\Big(-\partial_{+}P\partial_{-}P+M_{--}^{-1}\partial_{-}P\partial_{-}P$
$\hskip 96.73918pt-M_{--}(Q_{+}+J_{+}+M_{--}^{-1}(\partial_{-}P-J_{-}))^{2}$
$\hskip 96.73918pt+2(J_{+}-M_{--}^{-1}J_{-})\partial_{-}P-J_{+}J_{-}+M_{--}^{-1}J_{-}^{2}\Big)d^{2}x,$ (20)
with the notation $\partial_{\pm}P=\bar{e}_{\pm}{}^{\mu}\partial_{\mu}P$. To clean up this action let us shift $Q$ and $J$ are define
$Q^{\prime}_{+}=Q_{+}+J_{+}+M_{--}^{-1}(\partial_{-}P-J_{-}),$ (21)
and
${\cal J}_{+}=J_{+}-M_{--}^{-1}J_{-}.$ (22)
The action now takes the form
$S=S_{eff}+\frac{1}{2\pi}\int{\rm det}(\bar{e})M_{--}{Q^{\prime}_{+}}^{2}d^{2}x,$ (23)
where
$S_{eff}$ $=\frac{1}{2\pi}\int{\rm det}(\bar{e})\Big(\partial_{+}P\partial_{-}P-M_{--}^{-1}\partial_{-}P\partial_{-}P-2{\cal J}_{+}\partial_{-}P+{\cal J}_{+}J_{-}\Big)d^{2}x.$ (24)
Thus we see that the action splits into two decoupled parts parts: a contribution involving the scalar $P$ alone, coupled to sources, together with a simple quadratic term in the field $Q^{\prime}_{+}$. In particular the action for $Q^{\prime}_{+}$ involves no derivatives, so that $Q^{\prime}_{+}$ is an auxiliary field and so can be integrated out to give a result that is independent of $P$ and the sources.
2.3 Evaluating $M_{--}$
[…]
$g_{\mu\nu}=\bar{e}_{\mu}{}^{a}\bar{e}_{\nu}{}^{b}G_{ab}.$ (25)
We see that ${\det G}=-{\det g}/\det\bar{g}$ and this is non-vanishing so we can introduce the inverse $G^{ab}$. This allows us to write
$(\star F)_{a}$ $=\bar{e}_{a}{}^{\mu}(\star F)_{\mu}$
$=\sqrt{-{\det G}}\epsilon_{ab}G^{bc}F_{c}.$ (26)
From here we find
$\star F_{+}=\sqrt{-{\det G}}G^{-+}F_{+}+\sqrt{-G}G^{--}F_{-}$
$\star F_{-}=-\sqrt{-{\det G}}G^{+-}F_{-}-\sqrt{-G}G^{++}F_{+}.$ (27)
Thus if $\star F=F$ then
$F_{+}=\sqrt{-{\det G}}G^{-+}F_{+}+\sqrt{-{\det G}}G^{--}F_{-}$
$F_{-}=-\sqrt{-{\det G}}G^{+-}F_{-}-\sqrt{-{\det G}}G^{++}F_{+},$ (28)
from which we find
$F_{-}=\frac{1-\sqrt{-{\det G}}G^{-+}}{\sqrt{-{\det G}}G^{--}}F_{+},$ (29)
and
$F_{-}=-\frac{\sqrt{-{\det G}}G^{++}}{1+\sqrt{-{\det G}}G^{+-}}F_{+}.$ (30)
These two expressions are equal since $G^{-+}=G^{+-}$ and $G^{++}G^{--}-G^{-+}G^{+-}={\det G}^{-1}$.
In our case we have $F=Q+{\cal M}(Q)$ so that $F_{+}=Q_{+}$ and $F_{-}=M_{--}Q_{+}$. Therefore demanding that $F$ is self-dual with respect to $\star$ we read off that
$M_{--}=\frac{1-\sqrt{-{\det G}}G^{-+}}{\sqrt{-{\det G}}G^{--}}=-\frac{\sqrt{-{\det G}}G^{++}}{1+\sqrt{-{\det G}}G^{+-}}.$ (31)
Note that for a two-by-two matrix we have
$G^{ab}=\frac{1}{{\det G}}\left(\begin{array}[]{cc}G_{--}&-G_{-+}\\ -G_{+-}&G_{++}\end{array}\right),$ (34)
so
$M_{--}=-\frac{\sqrt{-{\det G}}-G_{+-}}{G_{++}}=\frac{G_{--}}{\sqrt{-{\det G}}+G_{-+}}.$ (35)
To continue, we choose a parameterisation of the two-dimensional metric $\bar{g}$ and give express $M_{--}$ in terms of this parameterisation. Without loss of generality, we can parameterise the metric in terms of three functions $\sigma(x),\rho(x),\tilde{\rho}(x)$ as
$\bar{g}=e^{2\sigma}(dx+\rho dy)(dx+\tilde{\rho}dy).$ (36)
A choice of zweibein is
$\bar{e}^{+}=\frac{1}{\sqrt{2}}e^{\sigma}(dx+\rho dy),\qquad\bar{e}^{-}=\frac{1}{\sqrt{2}}e^{\sigma}(dx+\tilde{\rho}dy).$ (37)
Then $\det\bar{g}=-(\det\bar{e})^{2}$ with
$\det\bar{e}=\frac{1}{2}e^{2\sigma}(\tilde{\rho}-\rho).$ (38)
Similarly for the physical metric we write
$g=e^{2\omega}(dx+\tau dy)(dx+\tilde{\tau}dy),$ (39)
for other functions $\omega(x),\tau(x),\tilde{\tau}(x)$. In this case we take
$\bar{e}^{+}=\frac{1}{\sqrt{2}}e^{\omega}(dx+\tau dy),\qquad\bar{e}^{-}=\frac{1}{\sqrt{2}}e^{\omega}(dx+\tilde{\tau}dy),$ (40)
and find $\det g=-(\det e)^{2}$ with
$\det e=\frac{1}{2}e^{2\omega}(\tilde{\tau}-\tau).$ (41)
A short computation shows that
[…]
and hence we obtain
$M_{--}$ $=\frac{(\rho-\tau)(\rho-\tilde{\tau})}{\tfrac{1}{2}(\rho-\tilde{\rho})(\tau-\tilde{\tau})-\tau\tilde{\tau}-\rho\tilde{\rho}+\tfrac{1}{2}(\tau+\tilde{\tau})(\rho+\tilde{\rho})}$
$=-\frac{\rho-\tau}{\tilde{\rho}-\tau}.$ (45)
In the special case $\rho=-\tilde{\rho}=1$ we have
$M_{--}$ $=\frac{(1-\tau)}{(1+\tau)}.$ (46)
[…]
Let us make some comments about $M_{--}$ in special cases. First we note that when $\rho=\tau$ we find $M_{--}=0$. This arises as the one form $dx+\rho dy=dx+\tau dy$ is then self-dual with respect to both $g$ and $\bar{g}$. If we also impose $\tilde{\rho}=\tilde{\tau}$, then $g$ is conformal to $\bar{g}$. Since the self-duality condition is conformally invariant we expect to find $M_{--}=0$ in this case. However, we also find $g$ is conformal to $\bar{g}$ if we take $\tilde{\rho}=\tau$ and $\tilde{\tau}=\rho$ but this corresponds to $M_{--}\to\infty$. The reason for this is that while $dx+\rho dy$ is always taken to be self-dual with respect to $\bar{g}$, $dx+\tilde{\tau}dy=dx+\rho dy$ is anti-self-dual with respect to $g$. Thus although the two metrics are conformal to each other the notion of self-duality has been flipped. As discussed in the introduction, we choose $\bar{g}$ so that $M_{--}$ is finite and non-zero everywhere, so for given $\tau,\tilde{\tau}$ we choose $\rho,\tilde{\rho}$ to exclude the above cases.
Note that the classical theory has two Weyl symmetries, one for each metric, under
$g_{\mu\nu}\to\Omega^{2}(x)g_{\mu\nu},\qquad\bar{g}_{\mu\nu}\to\bar{\Omega}^{2}(x)\bar{g}_{\mu\nu}$ (47)
as can be seen from the fact that $M$ doesn’t depend on $\omega$ or $\sigma$.
2.4 Evaluating $S_{eff}$
We now discuss further the effective action (24) that we found for $P$. It is helpful to write it in a more familiar form as
$S_{eff}$ $=\frac{1}{2\pi}\int\sqrt{-\det h}\Big(\frac{1}{2}h^{\mu\nu}\partial_{\mu}P\partial_{\nu}P-2{\cal J}^{\mu}\partial_{\mu}P+{\cal J}^{\mu}J_{\mu}\Big)d^{2}x,$ (48)
where
[…]
and
${\cal J}^{\mu}=\frac{\det\bar{e}}{\sqrt{-\det h}}\bar{e}_{-}{}^{\mu}{\cal J}_{+}.$ (50)
Note $h_{\mu\nu}$ is only determined up to a conformal factor that drops out of the action due to conformal invariance. A possible zweibein frame is
$h_{\mu\nu}=\tilde{e}_{\mu}{}^{a}\tilde{e}_{\nu}{}^{b}\eta_{ab},$ (51)
with
$\tilde{e}_{\mu}{}^{+}$ $=\bar{e}_{\mu}{}^{+}\qquad\tilde{e}_{\mu}{}^{-}=\bar{e}_{\mu}{}^{-}+M_{--}^{-1}\bar{e}_{\mu}{}^{+}$ (52)
$\tilde{e}_{+}{}^{\mu}$ $=\bar{e}_{+}{}^{\mu}-M_{--}^{-1}\bar{e}_{-}{}^{\mu}\qquad\tilde{e}_{-}{}^{\mu}=\bar{e}_{-}{}^{\mu}$. (53)
Its easy to see that $\det\tilde{e}=\det\bar{e}$ and hence $\sqrt{-\det h}=\det\bar{e}$ and ${\cal J}^{\mu}=\bar{e}_{-}{}^{\mu}{\cal J}_{+}=\tilde{e}_{-}{}^{\mu}{\cal J}_{+}$. In particular. given the metric parameterizations above we find
$\tilde{e}^{+}$ $=\frac{1}{\sqrt{2}}e^{\sigma}(dx+\rho dy)$ (54)
$\tilde{e}^{-}$ $=\frac{1}{\sqrt{2}}e^{\sigma}\frac{\rho-\tilde{\rho}}{\rho-\tau}\left(dx+\tau dy\right).$ (55)
By a local Lorentz transformation
$\tilde{e}^{\prime+}=\Lambda\tilde{e}^{+}\qquad\tilde{e}^{\prime-}=\Lambda^{-1}\tilde{e}^{-}$, (56)
with $\Lambda=\sqrt{(\rho-\tau)/(\rho-\tilde{\rho})}$, we can also find something more symmetric:
$\tilde{e}^{\prime+}$ $=\frac{1}{\sqrt{2}}e^{\sigma^{\prime}}(dx+\rho dy)$ (57)
$\tilde{e}^{\prime-}$ $=\frac{1}{\sqrt{2}}e^{\sigma^{\prime}}\left(dx+\tau dy\right).$ (58)
For either choice of the zweibein the metric $h$ can be taken to be
$h_{\mu\nu}dx^{\mu}dx^{\nu}$ $=e^{2\sigma^{\prime}}(dx+\rho dy)\left(dx+\tau dy\right),$ (59)
where we have introduced $\sigma^{\prime}=\sigma+\tfrac{1}{2}\ln((\rho-\tilde{\rho})/(\rho-\tau))$. However, we recall that the metric $h$ is only defined up to a conformal factor so we can choose $\sigma^{\prime}$ to be anything we like.
3 The Quantum Partition Function
3.1 The Path Integral And A Complex Metric
Let us now turn our attention to the quantum theory. We start with the path integral
$Z[J]$ $=\int[dP][dQ_{+}]e^{iS[P,Q_{+},J]}$
$=\int[dP][dQ^{\prime}_{+}]e^{iS_{eff}[P,J]+\frac{i}{2\pi}\int{\det}(\bar{e})M_{--}(Q^{\prime}_{+})^{2}}.$ (60)
Our first task is to perform the path integral over $Q^{\prime}_{+}$, which is a simple Gaussian giving formally an infinite factor $\prod_{x}({\det}(\bar{e})M_{--})^{-1/2}(x)$. In Andriolo:2021gen it was argued that this can be regularised to give a finite number that does not depend on ${\det}(\bar{e})M_{--}$. Dropping this irrelevant multiplicative factor, we find
$Z[J]\sim\int[dP]e^{iS_{eff}[P,J]}.$ (61)
In Andriolo:2021gen the path integral was made convergent by taking the metric $g$ to have Euclidean signature. Here, we wish to consider the path integral for the case in which both metrics $\bar{g}$ and $g$ have Euclidean signature. As discussed in the introduction, we actually consider the more general case in which $\bar{g}$ and $g$ are complex metrics, with $\rho,\tilde{\rho},\tau$ and $\tilde{\tau}$ independent complex variables. We then find Riemannian metrics $\bar{g}$ and $g$ in the case that $\tilde{\rho}=\rho^{*}$ and $\tilde{\tau}=\tau^{*}$ respectively, but we can also be more general. In what follows, we assume that spacetime is a compact Riemann surface; additional subtleties can arise on non-compact spaces.
To continue, we write $\sqrt{-h}=i\sqrt{h}$ and $S_{eff}=i\tilde{S}_{eff}$ with
$\tilde{S}_{eff}=\frac{1}{2\pi}\int\sqrt{h}\Big(\frac{1}{2}h^{\mu\nu}\partial_{\mu}P\partial_{\nu}P-2{\cal J}^{\mu}\partial_{\mu}P+{\cal J}^{\mu}J_{\mu}\Big)d^{2}x.$ (62)
so that now the path integral takes the form
$Z[\rho,\tau,J]$ $\sim\int[dP]e^{-\tilde{S}_{eff}[P,J]},$ (63)
which depends holomorphically on each of the two complex structures, $\rho$ and $\tau$. This is the same as a path integral for a non-chiral real scalar field but with a complex metric:
$h=e^{2\sigma^{\prime}}(dx+\rho dy)(dx+\tau dy).$ (64)
Here we see a complex metric combining the holomorphic sectors of two distinct Riemannian metrics arising naturally from the Sen action.
3.2 A Canonical Approach
It is instructive to construct the partition function from canonical quantization on a 2-torus. As usual, for real Riemannian metrics, the diffeomorphism symmetry can be used to bring the real metric $g$ on a 2-torus to standard form with $\tau$ constant, the complex structure modulus. As explained in Hull:2023dgp , there is a second diffeomorphism-like symmetry that can be used to bring the real metric $\bar{g}$ to standard form with $\rho$ a constant modulus. Here we set the sources to zero and take the functions $\rho$ and $\tau$ to be constants. As already mentioned we will extend our calculations to allow for $\rho,\tilde{\rho},\tau$ and $\tilde{\tau}$ to be general independent complex variables.
The metrics are then (using the symmetries to set the conformal factors to 1)
$g$ $=$ $(dx+\tau dy)(dx+\tilde{\tau}dy),$ (65)
$\bar{g}$ $=$ $(dx+\rho dy)(dx+\tilde{\rho}dy),$ (66)
$h$ $=$ $(dx+\tau dy)(dx+\rho dy).$ (67)
It is helpful to introduce the coordinates
$z=x+\tau y,\qquad\tilde{z}=x+\rho y.$ (68)
so that the metric is simply
$h=dzd\tilde{z}$. (69)
Note that for $\rho=-\tau=1$ we find that $h$ is the Minkowski metric with $y$ the time variable. Here we will be more general and allow for complex values of $\rho$ and $\tau$. The associated derivatives are
$\partial=-\frac{1}{\rho-\tau}\left(\partial_{y}-\rho\partial_{x}\right),\qquad\tilde{\partial}=\frac{1}{\rho-\tau}\left(\partial_{y}-\tau\partial_{x}\right),$ (70)
and these satisfy
$\partial z=\tilde{\partial}\tilde{z}=1,\qquad\partial\tilde{z}=\tilde{\partial}z=0.$ (71)
The equation of motion for $P$ becomes simply
$\partial\tilde{\partial}P=0,$ (72)
Assuming the coordinate $x$ is periodic: $x\cong x+2\pi l$, the general solution can be parameterised as
$P=P_{0}+\frac{1}{2l}a_{0}z-\frac{1}{2l}\tilde{a}_{0}\tilde{z}+\frac{i}{\sqrt{2}}\sum_{n\neq 0}\frac{a_{n}}{n}e^{-inz/l}-\frac{i}{\sqrt{2}}\sum_{n\neq 0}\frac{\tilde{a}_{n}}{n}e^{in\tilde{z}/l}.$ (73)
Note that the linear terms imply that $P(x+2\pi l,y)=P(x,y)+\pi(a_{0}-\tilde{a}_{0})$.
There are two cases to consider. First, if $P$ is a non-compact boson $P\in\mathbb{R}$, then we must have $a_{0}=\tilde{a}_{0}$ so setting
$a_{0}=p_{0},\qquad\tilde{a}_{0}=p_{0}$ (74)
we have
$P=P_{0}+\frac{1}{2l}p_{0}(\tau-\rho)y+\frac{i}{\sqrt{2}}\sum_{n\neq 0}\frac{a_{n}}{n}e^{-inz/l}-\frac{i}{\sqrt{2}}\sum_{n\neq 0}\frac{\tilde{a}_{n}}{n}e^{in\tilde{z}/l}$ (75)
with centre of mass position $P_{0}$ and continuous momentum proportional to $p_{0}$.
Secondly, if $P$ is periodic with $P\cong P+2\pi R$, then we must have $a_{0}-\tilde{a}_{0}=2mR$ for some constant $R$ and integer $m\in\mathbb{Z}$. To this end, we write
$a_{0}=p_{0}+mR,\qquad\tilde{a}_{0}=p_{0}-mR.$ (76)
Moreover, the momentum $p_{0}$ is discrete as the eigenvalues of $p_{0}$ are quantised: $p_{0}=n/R$ for integers $n\in\mathbb{Z}$. Then the states are labelled by two integers $m,n$, with $n/R$ the discrete momenta and $m$ being the winding number.
To quantize this system we treat $y$ as time and obtain the conjugate momentum
$\Pi$ $=\frac{i}{2\pi}\sqrt{h}h^{y\nu}\partial_{\nu}P$
$=\frac{1}{2\pi}(\partial-\tilde{\partial})P$
$=\frac{1}{2\pi l}\left(p_{0}+\frac{1}{\sqrt{2}}\sum_{n\neq 0}a_{n}e^{-inz/l}-\frac{1}{\sqrt{2}}\sum_{n\neq 0}{\tilde{a}_{n}}e^{in\tilde{z}/l}\right).$ (77)
To reproduce the quantization condition
$[P(x,y),\Pi(x^{\prime},y)]=i\delta(x-x^{\prime}),$ (78)
we find the non-zero commutation relations
$[a_{n},a_{m}]=n\delta_{m,-n},\qquad[\tilde{a}_{n},\tilde{a}_{m}]=n\delta_{m,-n},\qquad[P_{0},p_{0}]=i.$ (79)
The Hamiltonian is
$H$ $=\int\Pi\partial_{y}P-{\cal L}dx$
$=\frac{1}{2\pi}\int\left(\rho(\tilde{\partial}P)^{2}-\tau(\partial P)^{2}\right)dx$
$=\frac{\rho}{l}H_{\rho}-\frac{\tau}{l}H_{\tau},$ (80)
where
$H_{\tau}$ $=\frac{1}{4}(p_{0}+mR)^{2}+\sum_{n>0}a_{-n}a_{n}-\frac{1}{24}$
$H_{\rho}$ $=\frac{1}{4}(p_{0}-mR)^{2}+\sum_{n>0}\tilde{a}_{-n}\tilde{a}_{n}-\frac{1}{24},$ (81)
and as usual the 1/24-1/24 arises from normal ordering.
As is familiar in two dimensions, the left and right sectors decouple so that the total Hilbert space is a direct product ${\cal H}={\cal H}_{\tau}\otimes{\cal H}_{\rho}$. Here this corresponds to the decoupling of the two chiral bosons, as first observed in Sen:2019qit . Since the Hilbert space factorises the partition function takes the form
$Z(\beta)$ $={\rm Tr}_{{\cal H}}\left(e^{-\beta H}\right)$
$={\rm Tr}_{{\cal H}_{\tau}\otimes{\cal H}_{\rho}}\left(e^{\beta l^{-1}\tau H_{\tau}}e^{-\beta l^{-1}\rho H_{\rho}}\right)$
$={\rm Tr}_{{\cal H}_{\tau}\otimes{\cal H}_{\rho}}\left(q^{H_{\tau}}\tilde{q}^{H_{\rho}}\right),$ (82)
where $q=e^{\beta l^{-1}\tau}$ and $\tilde{q}=e^{-\beta l^{-1}\rho}$. This can be evaluated in the usual way. Summing over the oscillators gives
$Z(\beta)=\frac{1}{\eta(q)\eta(\tilde{q})}{\rm Tr}_{zero-modes}\left(q^{H_{\tau}}\tilde{q}^{H_{\rho}}\right)$ (83)
and it remains to take the trace of $q^{H_{\tau}}\tilde{q}^{H_{\rho}}$ over the zero-modes. If $P$ is a non-compact boson, then there are no winding numbers ($m=0$) and the momentum $p_{0}$ is a continuous variable that we integrate over to find
$Z(\beta)\sim\frac{1}{\sqrt{\rho-\tau}}\frac{1}{\eta(q)\eta(\tilde{q})}$. (84)
On the other hand, if $P$ takes values in a circle of radius $S$, then the eigenvalues of $p_{0}$ are quantised: $p_{0}=n/R$. Summing over $n$ and $m$ leads to
$Z(\beta)\sim\frac{1}{\eta(q)\eta(\tilde{q})}\sum_{m,n}q^{\frac{1}{4}(n/R+mR)^{2}}\tilde{q}^{\frac{1}{4}(n/R-mR)^{2}}.$ (85)
We will find agreement with the path integral calculation below by taking $\beta=2\pi il$, as expected.
3.3 Path Integral On The Torus
Our next task is to provide an explicit calculation of the partition function by evaluating the path integral on a torus. As above we take
$h=(dx+\rho dy)(dx+\tau dy),$ (86)
with $\rho$ and $\tau$ constant but complex and unrelated. We assume that $x,y\in[0,2\pi l]$ with periodic boundary conditions (except for the zero-modes). In this section we will closely follow the calculations in sections 10.2 and 10.4 of DiFrancesco:1997nk . The set-up discussed in DiFrancesco:1997nk arises in our analysis on setting $\rho=\tau^{*}$, but here we will keep $\rho$ and $\tau$ unrelated. A related calculation for the Sen action with $\rho=1$ was performed in Andriolo:2021gen and also Lambert:2023qgs .
First we split $P=P_{z.m.}+P_{o}$ where $P_{o}$ represents the non-zero oscillator modes which can be expanded in a basis of eigenmodes
$P_{o}=\sum_{n,m}c_{n,m}\phi_{n,m},$ (87)
where
$-\frac{1}{\sqrt{h}}\partial_{\mu}\left(\sqrt{h}h^{\mu\nu}\partial_{\nu}\phi_{n,m}\right)=\lambda_{n,m}\phi_{n,m},$ (88)
where $\sqrt{h}=\sqrt{det(h_{\mu\nu})}$ and we normalise
$\int d^{2}x\sqrt{h}\phi_{m,n}\phi_{-m^{\prime},-n^{\prime}}=\delta_{n,n^{\prime}}\delta_{m,m^{\prime}}.$ (89)
Taking a Fourier expansion we find
$\phi_{m,n}=Ne^{imx/l+iny/l},$ (90)
with $N^{-2}=4\pi^{2}l^{2}\sqrt{h}$ a normalization constant and
$\lambda_{m,n}=-\frac{4}{l^{2}}\frac{1}{(\rho-\tau)^{2}}(n-m\rho)(n-m\tau).$ (91)
The reality of $P$ is imposed by taking $c^{*}_{m,n}=c_{-m,-n}$.
The path integral leads to an infinite number of Gaussian integrals
$Z_{o}(\tau,\rho)$ $=\prod_{m,n}\int dc_{m,n}e^{-\frac{1}{4\pi}|c_{m,n}|^{2}\lambda_{m,n}+ic_{m,n}j_{-m,-n}},$ (92)
over the complex coefficients $c_{n,m}$ with $(m,n)\neq(0,0)$. Here $j_{m,n}$ arises from the sources:
${J}_{\mu}$ $=\sum_{m,n}{J}_{\mu m,n}e^{imx/l+iny/l}$
$j_{m,n}$ $=2\pi l\frac{\rho m-n}{\rho-\tau}(\tau{J}_{xm,n}-{J}_{ym,n}).$ (93)
The reality condition $c^{*}_{n,m}=c_{-n,-m}$ allows us to write this as
$Z_{o}(\tau,\rho)$ $=\prod_{(m,n)>0}\int dc_{m,n}dc^{*}_{m,n}e^{-\frac{1}{4\pi}c_{n,m}c^{*}_{n,m}\lambda_{m,n}+ic_{m,n}j_{-m,-n}+ic^{*}_{m,n}j_{m,n}}$
$=\prod_{(m,n)>0}\int dc^{\prime}_{m,n}d{c^{\prime}}^{*}_{m,n}e^{-\frac{1}{4\pi}c^{\prime}_{n,m}{c^{\prime}}^{*}_{n,m}\lambda_{n,m}+4\pi\lambda_{m,n}^{-1}j_{m,n}j_{-m,-n}},$ (94)
where we have restricted the product to be only over positive modes (for example only those pairs $(m,n)$ whose first non-zero entry is positive) and in the second line we have shifted the integration variables to remove the linear term. Thus we find
$Z_{o}=e^{-S_{n.l.}}\prod_{(m,n)>0}\frac{4\pi^{2}}{\lambda_{m,n}},$ (95)
where
$e^{-S_{n.l.}}=e^{4\pi\sum_{(m,n)>0}\lambda_{m,n}^{-1}j_{m,n}j_{-m,-n}}.$ (96)
However, since $\lambda_{n,m}$ is even in $(n,m)\to(-n,-m)$ we will find it easier to write this as
$Z_{o}(\tau,\rho)=e^{-S_{n.l.}}{\prod_{m,n}}^{\prime}\frac{2\pi}{\sqrt{\lambda_{m,n}}},$ (97)
where the product is over all $n,m\in\mathbb{Z}$, not both zero. We use the notation $\prod_{m,n}$$^{\prime}$ or $\sum_{m,n}$$^{\prime}$ to indicate a product or summation over all integers $m,n$ where not both are zero. As usual, to regulate this expression we define
$G(s)$ $=(4\pi^{2})^{s}{\sum_{m,n}}^{\prime}\lambda_{m,n}^{-s}$ (98)
$={(4\pi^{2}l^{2}\sqrt{h})^{s}}{\sum}^{\prime}_{m,n}\frac{1}{(n-m\rho)^{s}(n-m\tau)^{s}},$ (99)
and identify
$Z_{o}(\tau,\rho)=e^{-S_{n.l.}}e^{\frac{1}{2}G^{\prime}(0)}.$ (100)
Next, we write
$G(s)$ $={(4\pi^{2}l^{2}\sqrt{h})^{s}}\tilde{G}(s),$ (101)
so that we then have
$Z_{o}=(4\pi^{2}l^{2}\sqrt{h})^{\frac{1}{2}\tilde{G}(0)}e^{\frac{1}{2}\tilde{G}^{\prime}(0)}e^{-S_{n.l.}}.$ (102)
In other words, the coefficient $4\pi^{2}l^{2}\sqrt{h}$ simply provides an overall factor that multiplies the regulated infinite product.
Our next task is to compute $\tilde{G}(s)$, which exists for suitably large $s$, and with certain restrictions on $\rho$ and $\tau$, and then analytically continue it to $s=0$. In particular we find
$\tilde{G}(s)$ $={\sum_{n,m}}^{\prime}\frac{1}{(n-m\rho)^{s}(n-m\tau)^{s}}$
$=\sum_{n\neq 0}\frac{1}{n^{2s}}+\sum_{m\neq 0}\sum_{n}\frac{1}{(n-m\rho)^{s}(n-m\tau)^{s}}$
$=(1+(-1)^{2s})\zeta(2s)+\sum_{m\neq 0}\sum_{n}\frac{1}{(n-m\rho)^{s}(n-m\tau)^{s}}.$ (103)
To continue, we write
$\chi=\frac{1}{2}(\rho+\tau),\qquad\varphi=\frac{1}{2}(\rho-\tau),$ (104)
so that
$\tilde{G}(s)=(1+(-1)^{2s})\zeta(2s)+\sum_{m\neq 0}\sum_{n}\frac{1}{(n-m\chi-m\varphi)^{s}(n-m\chi+m\varphi)^{s}}.$ (105)
Here we notice that for fixed $m$ the sum over $n$ is periodic under $m\chi\to m\chi+1$ so we can make a Fourier expansion
$\tilde{G}(s)=(1+(-1)^{2s})\zeta(2s)+\sum_{m\neq 0}\sum_{n}\sum_{p}e^{2\pi imp\chi}\int_{0}^{1}dy\frac{e^{-2\pi ipy}}{(n-y-m\varphi)^{s}(n-y+m\varphi)^{s}}.$ (106)
Next, we note that the sum over $n$ converts the integral over $y$ from $[0,1]$ to $(-\infty,\infty)$:
$\tilde{G}(s)$ $=(1+(-1)^{2s})\zeta(2s)+\sum_{m\neq 0}\sum_{p}e^{2\pi imp\chi}\int_{-\infty}^{\infty}dy\frac{e^{-2\pi ipy}}{(-y-m\varphi)^{s}(-y+m\varphi)^{s}}$
$=(1+(-1)^{2s})\zeta(2s)+\sum_{m\neq 0}\sum_{p}e^{2\pi imp\chi}\int_{-\infty}^{\infty}dy\frac{e^{2\pi ipy}}{(y-m\varphi)^{s}(y+m\varphi)^{s}}$
$=(1+(-1)^{2s})\zeta(2s)+\sum_{m\neq 0}\sum_{p}e^{2\pi imp\chi}\int_{-\infty}^{\infty}dy\frac{e^{2\pi ipy}}{(y^{2}-m^{2}\varphi^{2})^{s}}.$ (107)
We now introduce a new integral:
$\tilde{G}(s)$ $=(1+(-1)^{2s})\zeta(2s)+\frac{1}{\Gamma(s)}\sum_{m\neq 0}\sum_{p}\int_{-\infty}^{\infty}dy\int_{0}^{\infty}\frac{dt}{t^{1-s}}e^{2\pi ip(m\chi+y)}e^{-t(y^{2}-m^{2}\varphi^{2})}.$ (108)
Note that convergence of the $t$ integral requires that ${\rm Re}(\varphi^{2})<0$. The integral over $y$ is Gaussian and can now be computed
$\tilde{G}(s)$ $=(1+(-1)^{2s})\zeta(2s)+\frac{\sqrt{\pi}}{\Gamma(s)}\sum_{m\neq 0}\sum_{p}\int_{0}^{\infty}\frac{dt}{t^{3/2-s}}e^{2\pi ipm\chi}e^{m^{2}t\varphi^{2}-\pi^{2}p^{2}/t}.$ (109)
We need to treat the $p=0$ term separately:
$\sum_{m\neq 0}\int_{0}^{\infty}\frac{dt}{t^{3/2-s}}e^{m^{2}t\varphi^{2}}$ $=\sum_{m\neq 0}\frac{\sqrt{\pi}}{\Gamma(s)}(-m^{2}\varphi^{2})^{1/2-s}$
$=2\frac{\sqrt{\pi}}{\Gamma(s)}(\sqrt{-\varphi^{2}})^{1-2s}\zeta(s-1/2),$ (110)
which again is valid if ${\rm Re}(\varphi^{2})<0$. Now we invoke the $\zeta$-function identity
$\pi^{-s^{\prime}/2}\Gamma(s^{\prime}/2)\zeta(s^{\prime}/2)=\pi^{(s^{\prime}-1)/2}\Gamma((1-s^{\prime})/2)\zeta(1-2s^{\prime}),$ (111)
to find something valid near $s=0$:
$\tilde{G}(s)$ $=(1+(-1)^{2s})\zeta(2s)+2\frac{\pi^{2s-1}}{\Gamma(s)}\Gamma(1-s)\zeta(2-2s)(\sqrt{-\varphi^{2}})^{1-2s}$
$\qquad+\frac{\sqrt{\pi}}{\Gamma(s)}\sum_{m\neq 0}\sum_{p\neq 0}\int_{0}^{\infty}\frac{dt}{t^{3/2-s}}e^{2\pi ipm\chi}e^{m^{2}t\varphi^{2}-\pi^{2}p^{2}/t}.$ (112)
To continue, we note that for small $s$, $\Gamma(s)\sim 1/s$ and so we can read off that
$\tilde{G}(0)=2\zeta(0)=-1.$ (113)
Furthermore, (noting that $\zeta^{\prime}(0)=-\tfrac{1}{2}\ln 2\pi$)
$\tilde{G}^{\prime}(0)$ $=(-\pi i-2\ln 2\pi)+2{\pi^{-1}}\Gamma(1)\zeta(2)\sqrt{-\varphi^{2}}$
$\qquad+{\sqrt{\pi}}\sum_{m\neq 0}\sum_{p\neq 0}e^{2\pi ipm\chi}\int_{0}^{\infty}\frac{dt}{t^{3/2}}e^{m^{2}t\varphi^{2}-\pi^{2}p^{2}/t}.$ (114)
This final integral can be computed (and exists for ${\rm Re}(\varphi^{2})<0$):
$\tilde{G}^{\prime}(0)$ $=(-\pi i-2\ln 2\pi)+\frac{\pi}{3}\sqrt{-\varphi^{2}}$
$\qquad+\sum_{m\neq 0}\sum_{p\neq 0}\frac{1}{|p|}e^{2\pi ipm\chi-2\pi|p||m|\sqrt{-\varphi^{2}}}$
$\sim\frac{\pi}{3}\sqrt{-\varphi^{2}}+2\sum^{\infty}_{m=1}\sum^{\infty}_{p=1}\frac{1}{p}e^{2\pi ipm\chi-2\pi pm\sqrt{-\varphi^{2}}}+\sum^{\infty}_{m=1}\sum^{\infty}_{p=1}\frac{1}{p}e^{-2\pi ipm\chi-2\pi pm\sqrt{-\varphi^{2}}}.$ (115)
The first term just gives a constant multiplicative factor and we have dropped it in the second line. If we introduce
$q=e^{2\pi i\chi-2\pi\sqrt{-\varphi^{2}}},\qquad\tilde{q}=e^{-2\pi i\chi-2\pi\sqrt{-\varphi^{2}}},$ (116)
then
$\tilde{G}^{\prime}(0)$ $\sim\frac{\pi}{3}\sqrt{-\varphi^{2}}+2\sum^{\infty}_{m=1}\sum^{\infty}_{p=1}\frac{1}{p}q^{pm}+\sum^{\infty}_{m=1}\sum^{\infty}_{p=1}\frac{1}{p}\tilde{q}^{pm}$
$\sim\frac{\pi}{3}\sqrt{-\varphi^{2}}-2\sum^{\infty}_{m=1}\ln(1-q^{m})-2\sum^{\infty}_{m=1}\ln(1-\tilde{q}^{m}).$ (117)
Let us recall that $\chi=(\rho+\tau)/2$ and $\varphi=(\rho-\tau)/2$. If we take
$\sqrt{-\varphi^{2}}=i(\rho-\tau)/2,$ (118)
then we find
$q=e^{2\pi i\tau},\qquad\tilde{q}=e^{-2\pi i\rho}.$ (119)
Putting all the pieces together, we find
$Z_{o}(\tau,\rho)\sim\frac{1}{\sqrt{\rho-\tau}}\frac{1}{\eta(q)}\frac{1}{\eta(\tilde{q})}e^{-S_{n.l.}},$ (120)
where
$\eta(q)=q^{1/24}\prod_{m=1}^{\infty}(1-q^{m}),$ (121)
and $(\rho-\tau)^{-\frac{1}{2}}$ comes from the $(4\pi l^{2}\sqrt{h})^{\frac{1}{2}\tilde{G}(0)}=(4\pi l^{2}\sqrt{h})^{-\frac{1}{2}}$ prefactor in (102).
Next we look at the zero-mode contribution. Here we allow for $P$ not to be single valued but rather to take values on a circle with $P\cong P+2\pi R$ for some fixed $R$ so that $P$ can have winding modes. Then with winding numbers $m,n$ on the $x$ and $y$ directions we have
$P_{z.m.}=p_{0}+\frac{mRx}{l}+\frac{nRy}{l}.$ (122)
Such field configurations satisfy $h^{\mu\nu}\partial_{\mu}\partial_{\nu}P_{z.m.}=0$. Evaluating the action on them gives
$\tilde{S}_{z.m.}$ $=\frac{1}{2\pi}\int\sqrt{h}\left(\frac{1}{2}h^{\mu\nu}\partial_{\mu}P_{z.m.}\partial_{\nu}P_{z.m.}-2{\cal J}^{\mu}_{0,0}\partial_{\mu}P_{z.m.}\right)d^{2}x$
$=\frac{\pi R^{2}}{\sqrt{h}}(n-\rho m)(n-m\tau)-\frac{4\pi lR}{\rho-\tau}(\tau J_{x0,0}-J_{y0,0})(n-\rho m).$ (123)
Noting that $\sqrt{h}=i(\rho-\tau)/2$, this leads to
$Z_{z.m.}(\tau,\rho)=\sum_{m,n}e^{-\frac{2\pi iR^{2}}{{\rho-\tau}}(n-\rho m)(n-m\tau)-\frac{4\pi lR}{{\rho-\tau}}(\tau J_{x0,0}-J_{y0,0})(n-\rho m)}.$ (124)
Next, we perform a Poisson re-summation over $n$:
$\sum_{n}e^{-\pi An^{2}+Bn}=\frac{1}{\sqrt{A}}\sum_{k}e^{-\frac{\pi}{A}\left(k+\frac{B}{2\pi i}\right)^{2}}.$ (125)
This requires that ${\rm Re}A>0$. Using $A=2iR^{2}/(\rho-\tau)$ and $B=\pi R^{2}(\rho+\tau)m/\sqrt{h}-4\pi lR(\tau J_{x0,0}-J_{y0,0})/(\rho-\tau)$, we get
$Z_{z.m.}(\tau,\rho)$ $=\frac{\sqrt{\rho-\tau}}{\sqrt{2i}R}\sum_{m,k}e^{\frac{2\pi iR^{2}}{{\rho-\tau}}\rho\tau m^{2}-\frac{\pi i({\rho-\tau})}{2R^{2}}k^{2}+\pi i(\rho+\tau)mk-\frac{\pi iR^{2}m^{2}}{2(\rho-\tau)}(\rho+\tau)^{2}}$
$\hskip 56.9055pt\times e^{2\pi l\left(ik/R-2i(\rho+\tau)mR/(\rho-\tau)+4imR\rho/(\rho-\tau)\right)(\tau J_{x0,0}-J_{y0,0})-\frac{2\pi il^{2}}{\rho-\tau}(\tau J_{x0,0}-J_{y0,0})^{2}}$
$=\frac{\sqrt{\rho-\tau}}{\sqrt{2i}R}\sum_{m,k}e^{-\frac{\pi iR^{2}}{2}(\rho-\tau)m^{2}-\frac{\pi i(\rho-\tau)}{R^{2}}k^{2}+\pi i(\rho+\tau)mk}$
$\hskip 56.9055pt\times e^{2\pi il(mR+k/R)(\tau J_{x0,0}-J_{y0,0})-\frac{2\pi il^{2}}{\rho-\tau}(\tau J_{x0,0}-J_{y0,0})^{2}}$
$=\frac{\sqrt{\rho-\tau}}{\sqrt{2i}R}e^{-\frac{2\pi i}{\rho-\tau}(\tau J_{x0,0}-J_{y0,0})^{2}}$
$\hskip 28.45274pt\times\sum_{m,k}e^{-2\pi il(mR+k/R)J_{y0,0}}q^{\frac{1}{4}(mR+k/R)^{2}+(mR+k/R)lJ_{x0,0}}\tilde{q}^{\frac{1}{4}(mR-k/R)^{2}}.$ (126)
Putting everything together gives:
$Z(\tau,\rho)$ $\sim\frac{1}{\eta(q)\eta(\tilde{q})}e^{-S_{n.l.}}e^{\frac{i}{2\pi}\int{\cal J}\wedge\tilde{\star}J}e^{-\frac{2\pi i}{\rho-\tau}(\tau J_{x0,0}-J_{y0,0})^{2}}$
$\hskip 28.45274pt\times\sum_{m,k}e^{-2\pi i(mR+k/R)J_{y0,0}}q^{\frac{1}{4}(mR+k/R)^{2}+(mR+k/R)lJ_{x0,0}}\tilde{q}^{\frac{1}{4}(mR-k/R)^{2}}.$ (127)
However, we can further simply this by separating out the zero modes in ${\cal J}\wedge\tilde{\star}J$
$e^{\frac{i}{2\pi}\int{\cal J}\wedge\tilde{\star}J}=$ $e^{\frac{i}{2\pi}\int{\cal J}_{o}\wedge\tilde{\star}J_{o}}e^{\frac{2\pi il^{2}}{\rho-\tau}(\tau J_{x0,0}-J_{y0,0})(\rho J_{x0,0}-J_{y0,0})},$ (128)
and hence we find
$Z(\tau,\rho)$ $\sim\frac{1}{\eta(q)\eta(\tilde{q})}e^{-S_{n.l.}}e^{\frac{i}{2\pi}\int{\cal J}_{o}\wedge\tilde{\star}J_{o}}e^{2\pi il^{2}(\tau J_{x0,0}-J_{y0,0})J_{x0,0}}$
$\hskip 28.45274pt\times\sum_{m,k}e^{-2\pi il(mR+k/R)J_{y0,0}}q^{\frac{1}{4}(mR+k/R)^{2}+(mR+k/R)lJ_{x0,0}}\tilde{q}^{\frac{1}{4}(mR-k/R)^{2}}.$ (129)
Note that the non-zero modes of the sources only appear in the first two exponential terms. It is interesting to observe that the source’s zero-modes only appear coupled to $\tau$ and not $\rho$. In the absence of sources we find more simply
$Z(\tau,\rho)$ $\sim\frac{1}{\eta(q)\eta(\tilde{q})}\sum_{m,k}q^{\frac{1}{4}(mR+k/R)^{2}}\tilde{q}^{\frac{1}{4}(mR-k/R)^{2}},$ (130)
which is of the form given in (4) that we discussed in the introduction and also agrees with the partition function (85) calculated by canonical quantisation (taking $\beta=2\pi il$).
3.4 Convergence
In the analysis above we have proceeded formally and assumed that the integrals we discussed are convergent where necessary. Let us now consider more carefully the criteria for their convergence. In full generality, we have a four-complex dimensional space of parameters $\rho,\tilde{\rho},\tau,\tilde{\tau}$. To simplify our analysis in this subsection, let us consider the case in which $\tilde{\rho}=\rho^{*}$ and $\tilde{\tau}=\tau^{*}$ which reduces our parameters to a two-dimensional complex subspace corresponding to the case where both $g$ and $\bar{g}$ are Riemannian.
We recall that our calculation consisted of two path integrals: one over $Q_{+}^{\prime}$, which leads to an effective action involving $P$, and then we must perform a path integral over $P$. We start with looking at when the integration over $Q^{\prime}_{+}$ is valid. In particular, the condition that $Q^{\prime}_{+}$ integral is a standard Gaussian is that ${\rm Im}({\det}(\bar{e})M_{--})>0$. As discussed earlier, the integral remains divergent even in this case, but can be regularised. If we were to consider the case of Lorentzian-signature metrics $\bar{g}$ and $g$, where all metric components are real, then we would have simply ${\rm Im}({\det}(\bar{e})M_{--})=0$ and hence the integral would not be Gaussian. Turning to the Euclidean context where $\rho$ and $\tau$ are complex we find
${\rm Im}(\det\bar{e}M_{--})=e^{2\sigma}\frac{\rho_{2}}{|\rho-\tau^{*}|^{2}}\left((\rho_{1}-\tau_{1})^{2}+\tau_{2}^{2}-\rho_{2}^{2}\right),$ (131)
where $\rho=\rho_{1}+i\rho_{2}$ and $\tau=\tau_{1}+i\tau_{2}$. We take the conformal factor $e^{2\sigma}$ to be real and positive. If $\rho_{2}<0$ then we must have
$\rho_{2}<-\sqrt{\tau_{2}^{2}+(\rho_{1}-\tau_{1})^{2}}.$ (132)
Alternatively for $\rho_{2}>0$ we find
$\rho_{2}<\sqrt{\tau_{2}^{2}+(\rho_{1}-\tau_{1})^{2}}.$ (133)
Either way, there is a open set of values of $\rho$ and $\tau$ for which the $Q_{+}^{\prime}$ integral is convergent.
Next, let us examine the convergence of the path integral arising from the integration over $P$. Looking at $S_{eff}$, we see that we expect this requires that $\sqrt{h}h^{\mu\nu}\partial_{\mu}P\partial_{\nu}P$ be positive definite. Note that this is a slightly weaker than the so-called KSW condition Kontsevich:2021dmb ; Witten:2021nzp which requires that, in addition to the 0-form $P$, the standard action for a 1-form also converges. A computation shows that the eigenvalues of ${\rm Re}(\sqrt{h}h^{\mu\nu})$ are
$\lambda_{\pm}=\frac{1}{|\rho-\tau|^{2}}\left(X\pm\sqrt{X^{2}-Y}\right),$ (134)
where
$X$ $=(1+|\rho|^{2})\tau_{2}-(1+|\tau|^{2})\rho_{2}$
$Y$ $=2\tau_{2}^{2}\rho_{2}^{2}-(\rho_{2}^{2}+\tau_{2}^{2})\rho_{2}\tau_{2}-(\rho_{1}-\tau_{1})^{2}\rho_{2}\tau_{2}.$ (135)
Thus positivity of ${\rm Re}(\sqrt{h}h^{\mu\nu})$ requires that both $\lambda_{+}$ and $\lambda_{-}$ are positive and hence $X>0$ and $Y>0$. In full generality, this is a slightly complicated condition. However, we can restrict to the case $\rho_{2}<0$ and $\tau_{2}>0$ so that $X,Y>0$ and, from (132), we also require
$\rho_{2}^{2}-\tau_{2}^{2}>(\rho_{1}-\tau_{1})^{2}.$ (136)
On the other hand, we can also consider the explicit path-integral calculation over $P$ that we performed above for the case of a torus. There we found that the integrals converge if ${\rm Re}(\varphi^{2})<0$ (recall $\varphi=\frac{1}{2}(\rho-\tau)$) which is equivalent to
$(\rho_{2}-\tau_{2})^{2}>(\rho_{1}-\tau_{1})^{2}.$ (137)
By writing $\rho_{2}=-(1+\alpha)\tau_{2}$ one sees that (136) implies (137). There was also a condition ${\rm Re}A>0$ in the Poisson re-summation of the zero-modes which becomes simply $\tau_{2}-\rho_{2}>0$. Lastly, we find that convergence of the sum over $m$ in the $\eta$-functions requires that $\tau_{2}>0$ and $\rho_{2}<0$. Thus the explicit torus calculation requires $\tau_{2}>0$, $\rho_{2}<0$ in addition to (136). These were referred to as the “microscopic” conditions in BenettiGenolini:2025jwe and we see that, as found there, they are stronger than the condition $X,Y>0$ and (131), which allow for more general solutions.
[…]
4 Modular Properties And Anomalies
4.1 Modular Properties
Geometrically modular invariance arises from requiring invariance under large diffeomorphisms of the torus. However, it is well-known that the partition function for a single chiral boson is not modular invariant. This is sometimes used as an argument why a single chiral boson cannot have a satisfactory action formulation as one would expect the path integral to give a modular invariant expression on the torus. Therefore it is of interest to examine what happens for Sen’s action and the bi-metric action.
Before discussing our result and its generalisations, let us review the partition function obtained from the textbook canonical quantisation analysis on a spatial circle $x\cong x+2\pi$. First, we consider a non-chiral, non-compact boson where the mode expansion is $P=p_{0}+Ex^{0}+oscillators$. The partition function is
$Z(\tau,\tau^{*})$ $=\int dE\sum_{oscillators}q^{L_{0}}(q^{*})^{\tilde{L}_{0}}$
$\sim\frac{1}{\sqrt{\tau-\tau^{*}}}\frac{1}{\eta(q)\eta(q^{*})}.$ (138)
Here the factor of $({\tau-\tau^{*}})^{-1/2}$ arises from the integral over the continuous $E$ variable. The $\eta$-functions arise from sums over the left and right moving oscillators. The zero-mode $p_{0}$ does not contribute to the energy – this follows from time translational invariance – and only contributes an irrelevant divergent factor to the partition function. This is well-known to be modular invariant.
In the case of a periodic scalar $P\cong P+2\pi R$, the mode expansion is $P=p_{0}+nx^{0}/Rl+mRx^{1}/l+oscillators$, and both the winding and $E$ variables become discrete: $m,n\in\mathbb{Z}$. Here find
$Z(\tau,\tau^{*})$ $=\sum_{m,n}\sum_{oscillators}q^{L_{0}}(q^{*})^{\tilde{L}_{0}}$
$\sim\frac{1}{\eta(q)\eta(q^{*})}\sum_{m,n}q^{\frac{1}{4}(mR+n/R)^{2}}(q^{*})^{\frac{1}{4}(mR-n/R)^{2}}.$ (139)
Again this is modular invariant. This partition function can be written as
$Z(\tau,\tau^{*})\sim\frac{1}{\eta(q)\eta(q^{*})}\sum_{(p_{L},p_{R})\in\Gamma}q^{\frac{1}{2}p_{L}^{2}}(q^{*})^{\frac{1}{2}p_{R}^{2}}.$ (140)
where
$p_{L}=\frac{1}{\sqrt{2}}(n/R+mR),\qquad p_{R}=\frac{1}{\sqrt{2}}(n/R-mR),$ (141)
and the sum is over all points $(p_{L},p_{R})$ in the Narain lattice $\Gamma$. The modular invariance can be seen to follow from the fact that the Narain lattice is even and self-dual.
For rational values of $R^{2}$, the Narain lattice decomposes into sublattices $\Gamma=\Gamma_{L}\oplus\Gamma_{R}$ with $p_{L}\in\Gamma_{L}$ and $p_{R}\in\Gamma_{R}$. Then
$Z(\tau,\tau^{*})\sim\frac{1}{\eta(q)\eta(q^{*})}\sum_{p_{L}\in\Gamma_{L}}\sum_{p_{R}\in\Gamma_{R}}q^{\frac{1}{2}p_{L}^{2}}(q^{*})^{\frac{1}{2}p_{R}^{2}}.$ (142)
This then takes the form of a sum of holomorphically factorised terms:
$Z(\tau,\tau^{*})\sim\frac{1}{\eta(q)\eta(q^{*})}\sum_{a}|f_{a}(q)|^{2},$ (143)
for a finite set of generalised theta functions $f_{a}(q)$ (see e.g. Blumenhagen:2013fgp ). In general the functions $f_{a}(q)$ transform into each other under modular transformations, so that all are needed for modular invariance in general.
For example, for the self-dual radius $R=1$ we find
$Z(\tau,\tau^{*})\sim\frac{1}{\eta(q)\eta(q^{*})}\sum_{m,n}q^{\frac{1}{4}(m+n)^{2}}(q^{*})^{\frac{1}{4}(m-n)^{2}}.$ (144)
This can then be rewritten as a sum of holomorphically factorised terms:
$Z(\tau,\tau^{*})\sim\frac{1}{\eta(q)\eta(q^{*})}(|\theta_{3}(\tau)|^{2}+|\theta_{2}(\tau)|^{2}),$ (145)
where the theta-functions are
$\theta_{3}(\tau)=\sum_{m\in\mathbb{Z}}q^{m^{2}},\qquad\theta_{2}(\tau)=\sum_{m\in\mathbb{Z}+\frac{1}{2}}q^{m^{2}}.$ (146)
and these transform into each other under modular transformations.
For rational $R^{2}$ there are states with $p_{R}=0$, i.e. there are values of momentum and winding numbers $m,n$ for which mRn/R=0mR-n/R=0. Restricting to these states defines a chiral boson theory with $p_{L}\in\Gamma_{L}$ and the partition function is
$Z(\tau)\sim\frac{1}{\eta(q)}\sum_{p_{L}\in\Gamma_{L}}q^{\frac{1}{2}p_{L}^{2}}.$ (147)
Here the self-duality condition implies that there is just one set of left-moving oscillators and hence a single $\eta$-function and the partition function is not modular invariant. For example, for the self-dual radius $R=1$ the states with $p_{R}=0$ have $m=n$ and $p_{L}=2m$ so that the partition function becomes
$Z(\tau)\sim\frac{1}{\eta(q)}\theta_{3}(q),$ (148)
and this is not modular invariant.
Note that if one tries to define a chiral boson theory when $R^{2}$ is not rational, the chiral constraint mRn/R=0mR-n/R=0 implies $m,n=0$ so that the only states are oscillators with no momentum or winding with $Z(\tau)\sim 1/\eta(q)$. This is not modular invariant either. Furthermore, in both these cases we have had to impose self-duality by hand, as an additional constraint that does not follow from the action.
We now look at the action of modular transformations on the partition function $Z(\tau,\rho)$ we have obtained above using the Sen formulation, which does arise directly from a Lagrangian formulation with no additional constraints. In the case $\rho=\tau^{*}$ then $Z(\tau,\tau^{*})$ agrees with the modular invariant partition function given in (139). This suggests that $Z(\tau,\rho)$ should be invariant under modular transformations in which both $\tau$ and $\rho$ transform. Indeed, it was shown in Hull:2023dgp the action (2.1) inherits two types of diffeomorphism-like symmetries that act separately on $g$ and $\bar{g}$. However, the ‘true’ diffeomorphisms, i.e. those that arise from coordinate transformations, correspond to the diagonal subgroup. Thus, looking at the global diffeomorphisms of the torus that generate modular transformations, this implies that $Z(\tau,\rho)$ should only be invariant under modular transformations where which both $\tau$ and $\rho$ transform. That is, the full modular group is $SL(2,\mathbb{Z})_{\tau}\times SL(2,\mathbb{Z})_{\rho}$ and $Z(\tau,\rho)$ should be invariant under the diagonal subgroup. Let us now confirm this.
First let us consider the case where $P$ is not periodic and so does not admit any winding modes. Here we find, setting the sources to zero for simplicity,
$Z(\tau,\rho)\sim Z_{o}(\tau,\rho)\sim\frac{1}{\sqrt{\tau-\rho}}\frac{1}{\eta(q)}\frac{1}{\eta(\tilde{q})}$. (149)
To examine its modular invariance we first note that
$\eta(\tau+1)=e^{\pi i/12}\eta(\tau),\qquad\eta(-1/\tau)=\sqrt{-i\tau}\eta(\tau).$ (150)
On the other hand we have defined $\eta(\rho)$ using $\tilde{q}=e^{-2\pi i\rho}$ which is valid in the lower half plane and so we have
$\eta(\rho+1)=e^{-\pi i/12}\eta(\rho),\qquad\eta(-1/\rho)=\sqrt{i\rho}\eta(\rho).$ (151)
As a result, for a non-periodic scalar the partition function does not factorise nor does it have any nice properties under general modular transformations of $\tau$ and $\rho$. However, it is invariant under diagonal modular transformations where $\tau$ and $\rho$ both transform in the same way.
Next we look at the case where $P$ is periodic: $P\cong P+2\pi R$. In addition to the non-zero mode contribution, we find the zero-mode contribution as given in (130). For rational values of $R$, the partition function takes the form
[…]
For example, at the self-dual radius $R=1$ (145) is replaced by
$Z(\tau,\rho)\sim\frac{1}{\eta(q)\eta(\tilde{q})}(\theta_{3}(\tau)\theta_{3}(\rho)+\theta_{2}(\tau)\theta_{2}(\rho)).$ (153)
Again it is easy to see that the partition function will not have any simple transformations properties under $\tau\to\tau+1$ unless we also take $\rho\to\rho+1$ where it is invariant. Next taking both $\tau\to-1/\tau$ and $\rho\to-1/\rho$ we find that
$\frac{(n-\rho m)(n-m\tau)}{\rho-\tau}$ $\to\frac{(n+m/\rho)(n+m/\tau)}{1/\tau-1/\rho}$
$=\frac{(n\rho+m)(n\tau+m)}{\rho-\tau}.$ (154)
This simply corresponds to swapping the sum $(m,n)\to(n,-m)$ in (124) and hence the zero-mode contribution is invariant under diagonal modular transformations.
Lastly we note that if we include the source zero-modes the diagonal modular invariance is preserved they transform under $\tau\to\tau+1,\rho\to\rho+1$ as
$J_{x,0,0}\to J_{x,0,0}\qquad J_{y0,0}\to J_{y0,0}+J_{x0,0},$ (155)
and under $\tau\to-1/\tau$, $\rho\to-1/\rho$ as
$J_{x,0,0}\to-J_{y,0,0}\qquad J_{y,0,0}\to J_{x,0,0}.$ (156)
4.2 Gravitational Anomalies
A single left-moving chiral scalar also has a 2-dimensional gravitational anomaly. A consistent quantum theory can then only be found by combining it with other degrees of freedom. For example, the gravitational anomaly of a left-moving chiral scalar can be cancelled by adding a right-moving chiral scalar or two right-moving real chiral Fermions. However, such a theory of a right-moving chiral scalar plus left-moving degrees of freedom is not modular invariant in general.
The modular invariant partition function for a non-chiral scalar given above was
$Z(q,q^{*})\sim\frac{1}{\eta(q)\eta(q^{*})}\sum_{(p_{L},p_{R})\in\Gamma}q^{\frac{1}{2}p_{L}^{2}}(q^{*})^{\frac{1}{2}p_{R}^{2}},$ (157)
and is a sum the sum over all points $(p_{L},p_{R})$ in the Narain lattice $\Gamma$. For rational values of $R^{2}$, points in the Narain lattice with $p_{L}=0$ correspond to a right-moving chiral scalar and the ones with $p_{R}=0$ correspond to a left-moving chiral scalar. The partition function for a left and right chiral scalar given by restricting the sum to these states is not modular invariant. For example, for the self-dual radius $R=1$ the restriction to chiral scalars is
$Z(q,q^{*})\sim\frac{1}{\eta(q)\eta(q^{*})}|\theta_{3}(\tau)|^{2},$ (158)
which is not modular invariant. For the unrestricted scalar, the modular invariant partition function (145) has another term involving $\theta_{2}$ coming from the points in the Narain lattice where both $p_{L},p_{R}$ are non-zero.
As mentioned above, in the bi-metric chiral boson action, there are two gauge symmetries, one for which $g$ is the gauge field and one for which $\bar{g}$ is the gauge field, with the ‘true’ diffeomorphisms arising as a diagonal subgroup of these Hull:2023dgp . Anomalies arise for both these symmetries and these will be discussed more fully elsewhere (see Hull:2025bqo and future work).
The action for a left-moving chiral scalar (2.1) is invariant under both these symmetries, but the quantum theory has an anomaly in both. This can be viewed as the physical scalar $A$ coupling to $g$ giving rise to an anomaly in the $g$-symmetry and the shadow sector $C$ coupling to $\bar{g}$ giving rise to an anomaly in the $\bar{g}$-symmetry. However, if one adds a similar action for a right-moving chiral scalar the anomalies in both symmetries are cancelled, as this introduces a right-moving physical scalar $\tilde{A}$ coupling to $g$ and a right-moving shadow scalar $\tilde{C}$ coupling to $\bar{g}$. However, this theory is not invariant under either the $\tau$ or $\rho$ modular transformations for reasons similar to those described above. Such theories must then be coupled to other fields to achieve modular invariance. This will be explored in the following subsection.
[…]
5 Chiral Bosons On A Torus, Anomalies And Modular invariance
5.1 Scalars On A Torus
Consider first the standard theory of $n$ scalars taking values on an $n$-torus $T=\mathbb{R}^{n}/\Gamma$ for some lattice $\Gamma$, coupled to a world-sheet metric $g$. A constant $B$-field is incorporated by adding a topological term to the action that does not affect the equations of motion but which changes the definition of the canonical momentum. Then the momentum and winding vectors combine into a $2n$-vector $(p_{L}^{i},p_{R}^{i})$ taking values in the $2n$-dimensional Narain lattice $\Lambda$, where $p_{L}^{i}$ is the momentum for left-movers and $p_{R}^{i}$ is the momentum for right-movers ($i=1,\dots,n$). The partition function is then a straightforward generalisation of (175):
$Z(\tau,\tau^{*})\sim\frac{1}{\eta(q)^{n}\eta(q^{*})^{n}}\sum_{(p_{L},p_{R})\in\Lambda}q^{\frac{1}{2}p_{L}^{2}}(q^{*})^{\frac{1}{2}p_{R}^{2}}.$ (159)
This is modular invariant as the $2n$-dimensional Narain lattice $\Lambda$ is even and self-dual.
For special choices of the lattice and B-field (i.e. at special points in the Narain moduli space) the Narain lattice decomposes into sublattices $\Lambda=\Gamma_{L}\oplus\Gamma_{R}$ with $p_{L}\in\Gamma_{L}$ and $p_{R}\in\Gamma_{R}$. These special lattices are the rational lattices. For example, for $n=1$ the 2-dimensional lattice is rational when $R^{2}$ is rational. For such a rational torus, the partition function can be written as
$Z(\tau,\tau^{*})\sim\frac{1}{\eta(q)^{n}\eta(q^{*})^{n}}\sum_{p_{L}\in\Gamma_{L}}\sum_{p_{R}\in\Gamma_{R}}q^{\frac{1}{2}_{L}^{2}}(q^{*})^{\frac{1}{2}p_{R}^{2}}.$ (160)
Scalars on a rational torus define a rational conformal field theory, so that the partition function can be written in the form of a finite sum
$Z(\tau,\tau^{*})\sim\sum_{a,\bar{b}}C_{a\bar{b}}\chi_{a}(\tau)\chi_{\bar{b}}(\tau^{*})$ (161)
Here $C_{a\bar{b}}$ are constants, $a$ labels the left-moving primary fields and $\chi_{a}(\tau)$ is the character of the Verma module generated by the $a$’th primary field, while $\bar{b}$ labels the right-moving primary fields and $\chi_{\bar{b}}(\tau^{*})$ is the character of the Verma module generated by the $\bar{b}$’th primary field.
On a rational torus, one can restrict to $n$ chiral bosons with $p_{R}^{i}=0$ and $p_{L}\in\Gamma^{\prime}$ and the partition function is
$Z(\tau)\sim\frac{1}{\eta(q)^{n}}\sum_{p_{L}\in\Gamma_{L}}q^{\frac{1}{2}p_{L}^{2}}.$ (162)
This is not modular invariant in general.
5.2 The Action for Chiral Bosons on a Torus Target Space
The action (18) can be generalised to $n$ scalars $P^{i}$ and $n$ 1-forms $Q_{i+}$ (with $i=1,\dots n$) with a constant metric $G_{ij}$ and constant B-field $B_{ij}$. The action is (including sources $J_{i\pm}$)
$S$ $=-\frac{1}{2\pi}\int{\rm det}(\bar{e})\Big(-\partial_{+}P^{i}\partial_{-}P^{j}(G_{ij}+B_{ij})-2Q_{i+}\partial_{-}P^{i}-(Q_{+}+J_{+})_{i}(Q_{+}+J_{+})_{j}G^{ij}M_{--}$
$\hskip 85.35826pt+2G^{ij}(Q_{i+}J_{j-}+J_{i+}J_{j-})\Big)d^{2}x.$ (163)
We will take $G_{ij}=\delta_{ij}$ and take the $P^{i}$ to be coordinates on a torus given by $T=\mathbb{R}^{n}/\Gamma$ for some lattice $\Gamma$, with $P^{i}\sim P^{i}+V^{i}$ for any $V^{i}\in\Gamma$. By a change of coordinates, this could equivalently be written with scalars on a square lattice with $P^{i}\sim P^{i}+2\pi$ but with a different metric, $G_{ij}\neq\delta_{ij}$. As before, the term with $B_{ij}$ is a topological term that does not affect the equations of motion but which changes the definition of the canonical momentum.
The calculation of the partition function proceeds as before, with the result given by a sum over vectors in the Narain lattice $\Lambda$
$Z(\tau,\rho)\sim\frac{1}{\eta(q)^{n}\eta(\tilde{q})^{n}}\sum_{(p_{L},p_{R})\in\Lambda}q^{\frac{1}{2}p_{L}^{2}}(\tilde{q})^{\frac{1}{2}p_{R}^{2}}.$ (164)
This agrees with (159) on replacing $q^{*}$ with $\tilde{q}$ and agrees with (130) when $n=1$. At points in the Narain moduli space at which the Narain lattice is rational, the partition function takes the form
$Z(\tau,\rho)\sim\sum_{a,\bar{b}}C_{a\bar{b}}\chi_{a}(\tau)\chi_{\bar{b}}(\rho).$ (165)
5.3 Holomorphic Factorisation And Even Self-Dual Lattices
For a standard bosonic CFT the sum over oscillator modes gives a contribution to the partition function of the form
$Z_{o}(q,q^{*})\sim\frac{1}{\eta(q)^{n}\eta(q^{*})^{n}},$ (166)
which factorises into a function of $\tau$ times a function of $\tau^{*}$, but the inclusion of momenta and winding from the zero modes in general gives a function of both $\tau$ and $\tau^{*}$ that doesn’t factorise. For the special tori for which the CFT is rational, the partition function is a finite sum of terms, each of which factorises into a function of $\tau$ and a function of $\tau^{*}$.
For the Sen-type chiral boson action, the sum over oscillators gives a term which factorises into the product of a function of $\tau$ and a function of $\rho$, but now the momentum and winding contributions give a function of both $\tau$ and $\rho$ in general. Just as before, in the rational case, this can be written as a finite sum of terms, each of which factorises into a function of $\tau$ times a function of $\rho$.
For certain lattices, there can be points in the Narain moduli space for which there is a complete factorisation of the full partition function (including zero modes)
$Z(\tau,\tau^{*})\sim F(\tau)G(\tau^{*}),$ (167)
for some functions $F,G$ and such points will have an important role here. Such a complete factorisation arises at rational points in the Narain moduli space at which $\Lambda=\Gamma_{L}\oplus\Gamma_{R}$ with $\Gamma_{L}=\Gamma_{R}=\Delta$ where $\Delta$ is an even self-dual lattice. This can only occur if $n$ is a multiple of $8$. The partition function is then of the form
$Z(\tau,\tau^{*})\sim F(\tau)F^{*}(\tau^{*}).$ (168)
For example, for $n=8$ there is only one even self-dual lattice and in this case $\Delta$ is the root lattice of $E_{8}$. In this case the partition factorises as (168) where Lerche:1988np  [^1] Note that Lerche:1988np uses a convention where $\alpha^{\prime}=2$ but we have $\alpha^{\prime}=1$.
$F(\tau)=\frac{1}{\eta(q)^{8}}\Theta_{\Delta}(\tau),$ (169)
is the $\hat{E}_{8}$ character and
$\Theta_{\Delta}(\tau)=\sum_{p\in\Delta}q^{\frac{1}{2}p^{2}},$ (170)
is the theta-function for the $E_{8}$ lattice. The $E_{8}$ theta function can be written in terms of Jacobi theta functions as
$\Theta_{\Delta}(\tau)=\frac{1}{2}(\theta_{2}^{8}(\tau)+\theta_{3}^{8}(\tau)+\theta_{4}^{8}(\tau)).$ (171)
In these cases, the chiral partition function for chiral bosons can be taken to be
$Z(\tau)\sim F(\tau).$ (172)
Similarly for the corresponding Sen-type chiral boson theory the partition function is
$Z(\tau,\rho)\sim F(\tau)F^{*}(\rho).$ (173)
Next we consider a theory with $n$ chiral bosons with partition function $Z(\tau)\sim F(\tau)$ and embed it in a larger theory with other fields which couple to $g$ but not $\bar{g}$. For example we will consider the heterotic string in the next section. If the additional fields have a modular partition function $f(\tau,\tau^{*})$ then the total partition function
$Z(\tau,\tau^{*})\sim f(\tau,\tau^{*})F(\tau),$ (174)
is modular invariant. If the $n$ chiral bosons are replaced by the Sen-type chiral boson theory with partition function $Z(\tau,\rho)\sim F(\tau)F^{*}(\rho)$, then the partition function for the whole system is of the form
$Z(\tau,\rho)\sim f(\tau,\tau^{*})F(\tau)F^{*}(\rho),$ (175)
which is still invariant under $\tau$ modular transformations. On the other hand, in the general case in which $Z(\tau,\rho)$ is a sum of terms that each factorise, then modular invariance would be hard, if not impossible, to achieve.
Note that the partition function (175) may now have gravitational anomalies associated to $\bar{g}$. If $\bar{g}$ is treated as a fixed background field, then this may not be important. On the other hand, if we define a string-like theory in which we integrate over both $\bar{g}$ as well as the physical metric $g$, then $\rho$ modular invariance and no gravitational anomalies would be desirable. This could be achieved by adding further matter coupling to $\bar{g}$ but not $g$ with partition function $\tilde{f}(\rho,\rho^{*})$ such that $\tilde{f}(\rho,\rho^{*})F^{*}(\rho)$ is modular invariant and anomaly free. Then the total partition function
$Z(\tau,\rho)\sim f(\tau,\tau^{*})\tilde{f}(\rho,\rho^{*})F(\tau)F^{*}(\rho),$ (176)
would be invariant under separate $\tau$ and $\rho$ modular transformations. Let us now consider the more specific example of the heterotic string.
5.4 The Heterotic String
The heterotic string can be formulated with 16 chiral bosons on $\mathbb{R}^{16}/\Delta$ where $\Delta$ is a 16-dimensional even self-dual lattice. There are precisely two such lattices: the root lattice of $E_{8}\times E_{8}$ and the weight lattice of $Spin(32)/\mathbb{Z}_{2}$. These both have the partition function
$F(\tau)=\frac{1}{\eta(q)^{16}}\Theta_{\Delta}(\tau),$ (177)
where $\Theta_{\Delta}(\tau)$ is given by (170).
The remaining degrees if freedom consist of 10 non-chiral uncompactified scalars $X^{\mu}$, 10 right-moving fermions $\psi^{\mu}$, the $b,c$ ghosts and the right-moving superghosts $\beta,\gamma$. Then the total partition function is (174) where
$f(\tau,\tau^{*})=Z_{b}(\tau,\tau^{*})Z_{f}(\tau,\tau^{*}),$ (178)
where
$Z_{b}(\tau,\tau^{*})=\frac{1}{(\tau-\tau^{*})^{4}}\frac{1}{\eta(q)^{4}\eta(q^{*})^{4}},$ (179)
is the partition function for 8 non-chiral uncompactified transverse scalars and
$Z_{f}(\tau,\tau^{*})=\frac{1}{(\tau-\tau^{*})^{4}}(\theta_{3}^{4}(\tau^{*})-\theta_{2}^{4}(\tau^{*})-\theta_{4}^{4}(\tau^{*})),$ (180)
is the partition function for 8 transverse right-moving fermions, with the 3 terms coming from the sum over spin structures. [^2] This is the counting from the light-cone gauge formulation. In the covariant approach, $Z_{b}$ is the parition function for 10 scalars and the bosonic ghosts, while $Z_{b}$ is the parition function for 10 right-moving fermions and the superghosts. For either choice of lattice $\Delta$, the resulting partition function
$Z(\tau,\tau^{*})=Z_{b}(\tau,\tau^{*})Z_{f}(\tau,\tau^{*})F(\tau),$ (181)
is modular invariant.
Now we consider the same system of fields $X^{\mu},\psi^{\mu},b,c,\beta,\gamma$ coupling to the physical metric $g$ together with the Sen-type action for fields $Q_{i},P^{i}$ with $i=1,\dots,16$ coupled to both $g$ and $\bar{g}$, with the Narain lattice for the scalars $P^{i}$ chosen to be $\Lambda=\Delta\oplus\Delta$. The Sen-type action gives 16 left-moving scalars $A^{i}$ coupling to $g$ and 16 left-moving scalars $C^{i}$ coupling to $\bar{g}$. The quantisation gives the partition function
$Z(\tau,\tau^{*},\rho)=Z_{b}(\tau,\tau^{*})Z_{f}(\tau,\tau^{*})F(\tau)F(\rho),$ (182)
and this is invariant under the $\tau$ modular transformations, but not the $\rho$ modular transformations. This construction cancels the gravitational anomaly for the $g$-gauge symmetries but not for the $\bar{g}$-gauge symmetries.
Invariance under the $\rho$ modular transformations can be achieved by adding a 2nd set of fields $\tilde{X}^{\mu},\tilde{\psi}^{\mu},\tilde{b},\tilde{c},\tilde{\beta},\tilde{\gamma}$ coupling to $\bar{g}$ so that the partition function becomes
$Z(\tau,\tau^{*},\rho,\rho^{*})=[Z_{b}(\tau,\tau^{*})Z_{f}(\tau,\tau^{*})F(\tau)][Z_{b}(\rho,\rho^{*})Z_{f}(\rho,\rho^{*})F(\rho)],$ (183)
which is invariant under both $\tau$ and $\rho$ modular transformations. Then the system consists of a physical heterotic string $X^{\mu},\psi^{\mu},b,c,\beta,\gamma$ coupling to the physical metric $g$ (and the physical world-sheet gravitino $\chi$) together with a shadow heterotic string $\tilde{X}^{\mu},\tilde{\psi}^{\mu},\tilde{b},\tilde{c},\tilde{\beta},\tilde{\gamma},C^{i}$ coupling to $\bar{g}$ (and a shadow world-sheet gravitino $\tilde{\chi}$) with no coupling between the physical and shadow string theories, so that the shadow sector decouples and doesn’t affect any of the physics of the physical sector. Moreover, this construction cancels the gravitational anomaly for the $\bar{g}$-gauge symmetries, so that both the gravitational anomaly for the $g$-gauge symmetries and that for the $\bar{g}$-gauge symmetries are both anomaly-free.
Remarkably, a similar structure was found in Sen’s heterotic string field theory Sen:2015uaa and the new heterotic string field theory Hull:2025mtb , both of which are formulated in terms of two string fields $\Psi,\tilde{\Psi}$ with $\Psi$ representing all the physical states of the heterotic string while $\tilde{\Psi}$ gives rise to a second shadow heterotic string with the same spectrum but which decouples from the physical string.
6 Path Integral On A General Riemann Surface
In this section we extend the path integral discussion to higher genus Riemann surfaces. To this end, we write the effective action as
$\tilde{S}_{eff}$ $=-\frac{1}{2\pi}\int\Big(\frac{1}{2}dP\wedge\tilde{\star}dP-2{\cal J}\wedge\tilde{\star}dP+{\cal J}\wedge\tilde{\star}J\Big)$
$=-\frac{1}{2\pi}\int\Big(-\frac{1}{2}P\wedge\tilde{\star}\tilde{\Delta}P-2P\wedge d\tilde{\star}{\cal J}+{\cal J}\wedge\tilde{\star}J\Big).$ (184)
Here $\tilde{\star}$ is defined in analogy with the Hodge dual:
$\tilde{\star}V_{\mu}=\sqrt{h}\epsilon_{\mu\nu}h^{\nu\lambda}V_{\lambda},$ (185)
with $\tilde{\star}^{2}=-1$ when acting on odd-forms, $\tilde{\star}^{2}=1$ on even-forms, and
$\tilde{\Delta}P=\tilde{\star}d\tilde{\star}dP,$ (186)
is the Laplacian associated to the complex metric $h$. We can see here that the path integral takes the form
$Z[J]=Z_{z.m.}Z_{o}e^{\frac{i}{2\pi}\int{\cal J}\wedge\tilde{\star}\tilde{J}}$, (187)
where $Z_{z.m.}$ and $Z_{o}$ are the contributions arising from the zero-modes and non-zero-modes respectively.
To continue, let us first consider the non-zero modes. Here we can start from the second form of the expression in (6). We can shift the integral over $P$ to $P^{\prime}=P+B$ with $B$ determined by
$\tilde{\Delta}B=-2\tilde{\star}d\tilde{\star}{\cal J}.$ (188)
This is possible since, for the non-zero modes, $\tilde{\Delta}$ is invertible allowing us to obtain $B$. This removes the linear term in $P$ from the action. However, the price of doing this is to introduce a non-local term in $S_{eff}$ of the form
$S_{n.l.}=-\frac{1}{2\pi}\int dB\wedge\tilde{\star}{\cal J}.$ (189)
Thus we find
$Z$ $=Z_{z.m}\int[dP^{\prime}]_{o}e^{-\frac{1}{4\pi}\int P^{\prime}\wedge\tilde{\star}\tilde{\Delta}P^{\prime}}e^{\frac{1}{2\pi}\int{\cal J}\wedge\tilde{\star}J}e^{-S_{n.l}}$
$\sim\frac{1}{\sqrt{{\rm det}^{\prime}(\tilde{\Delta})}}Z_{z.m}e^{-S_{n.l}}e^{\frac{1}{2\pi}\int{\cal J}\wedge\tilde{\star}J},$ (190)
where the determinant only includes the non-zero modes of $\tilde{\Delta}$ and $Z_{z.m}$ represents the contributions from the zero-modes.
Choosing coordinates where $\rho$ and $\tau$ constant the operator $\tilde{\Delta}$ factorizes: $\tilde{\Delta}P=4\partial\tilde{\partial}P$ where, as above,
$\partial=-\frac{1}{\rho-\tau}\left(\partial_{y}-\rho\partial_{x}\right)\qquad\tilde{\partial}=\frac{1}{\rho-\tau}\left(\partial_{y}-\tau\partial_{x}\right).$ (191)
This means that the eigenvalues of the Laplacian take the following form:
$\lambda_{\vec{n}}=-\frac{4}{(\rho-\tau)^{2}}\lambda^{(\rho)}_{\vec{n}}\lambda^{(\tau)}_{\vec{n}}$, (192)
where $\vec{n}$ is a generic label for the eigenvalues and $\lambda_{\vec{n}}^{(\rho)}$ and $\lambda_{\vec{n}}^{(\tau)}$ only depend on $\rho$ and $\tau$ respectively. Formally the determinant is given by
${\rm det}^{\prime}(\tilde{\Delta})=\prod_{{\vec{n}}}\left(\frac{-2i}{\rho-\tau}\right)^{2}\lambda^{(\rho)}_{\vec{n}}\lambda^{(\tau)}_{\vec{n}}$. (193)
As before there will be a reality constraint on $P$ that relates the associated eigenfunctions and thereby restricts the product. Using a zeta-function regularization this product is defined as
${\rm det}^{\prime}(\tilde{\Delta})=e^{G^{\prime}(0)},$ (194)
where
$G(s)$ $=\left(\frac{-2i}{\rho-\tau}\right)^{2s}\sum_{\vec{n}}(\lambda^{(\rho)}_{\vec{n}})^{s}(\lambda^{(\tau)}_{\vec{n}})^{s}$
$\equiv\left(\frac{2i}{\rho-\tau}\right)^{2s}\tilde{G}(s).$ (195)
It follows that
$G^{\prime}(0)$ $=2\ln\left(\frac{-2i}{\rho-\tau}\right)\tilde{G}(0)+\sum_{\vec{n}}\ln\lambda^{(\rho)}_{\vec{n}}+\sum_{\vec{n}}\ln\lambda^{(\tau)}_{\vec{n}}$, (196)
and hence the determinant takes the form
${\rm det}^{\prime}(\tilde{\Delta})=\left(\frac{2i}{\rho-\tau}\right)^{2\tilde{G}(0)}D_{+}(\rho)D_{-}(\tau).$ (197)
In this way the path integral on a generic Riemann surface can be written as
$Z$ $\sim\left(\frac{-2i}{\rho-\tau}\right)^{-2\tilde{G}(0)}\frac{1}{\sqrt{D_{+}(\rho)D_{-}(\tau)}}Z_{z.m}e^{-S_{n.l}}e^{\frac{1}{2\pi}\int{\cal J}\wedge\tilde{\star}J}.$ (198)
To evaluate $Z_{z.m.}$ we recall that the zero-modes satisfy $\tilde{\Delta}P=0$. To obtain these we note that $dP$ is a one-form and hence can be written as $dP=\alpha_{+}+\alpha_{-}$ where $\tilde{\star}\alpha_{\pm}=\pm i\alpha_{\pm}$ and $d\alpha_{+}+d\alpha_{-}=0$. The condition $\tilde{\Delta}P=0$ now imposes that $id\alpha_{+}-id\alpha_{-}=0$. Thus we find that $d\alpha_{+}=d\alpha_{-}=0$ and $d\tilde{\star}\alpha_{+}=d\tilde{\star}\alpha_{-}=0$. In the case of a real Euclidean metric on a Riemann surface these correspond to the holomorhic and anti-holomorphic one-forms respectively and $\alpha_{+}=(\alpha_{-})^{*}$.
Furthermore, these can have non-zero periods over the 1-cycles. This in turn means that $P$ is not globally defined but is allowed to be multi-valued although the periods are restricted, for example to be integer. This means that the integration by parts that we used above is not valid.
To continue, we write $\alpha_{+}=\sum_{i}\hat{m}_{i}\alpha^{i}_{+}$ and $\alpha_{+}=\sum_{i}\hat{n}_{i}\alpha_{-}^{i}$ where the $i$ index labels 1-cycles of the Riemann surface. Using the Riemann bi-linear identity we find, starting from the first line of (6),
$Z_{z.m.}$ $={\rm exp}\left(\frac{1}{4\pi}\sum_{\hat{m}^{i},\hat{n}^{j}}\int\Big(i(\hat{m}_{i}\alpha^{i}_{+}+\hat{n}_{i}\alpha^{i}_{-})\wedge(\hat{m}_{j}\alpha^{j}_{+}-\hat{n}_{j}\alpha^{j}_{-})-2i{\cal J}\wedge(\hat{m}_{i}\alpha^{i}_{+}-\hat{n}_{i}\alpha^{i}_{-})\Big)\right)$
$={\rm exp}\Bigg(-\frac{i}{2\pi}\sum_{\hat{m}^{i},\hat{n}^{j}}\hat{m}_{i}\hat{n}_{j}\sum_{k}\Big(\oint_{A^{k}}\alpha_{+}^{i}\oint_{B_{k}}\alpha_{-}^{j}-\oint_{B_{k}}\alpha_{+}^{i}\oint_{A^{k}}\alpha_{-}^{j}\Big)$
$\qquad\qquad-\frac{i}{2\pi}\sum_{\hat{m}_{i}}\hat{m}_{i}\sum_{k}\Big(\oint_{A^{k}}{\cal J}\oint_{B_{k}}\alpha_{+}^{i}-\oint_{B_{k}}{\cal J}\oint_{A^{k}}\alpha_{+}^{i}\Big)$
$\qquad\qquad+\frac{i}{2\pi}\sum_{\hat{n}_{i}}\hat{n}_{i}\sum_{k}\Big(\oint_{A^{k}}{\cal J}\oint_{B_{k}}\alpha_{-}^{i}-\oint_{B_{k}}{\cal J}\oint_{A^{k}}\alpha_{-}^{i}\Big)\Bigg).$ (199)
To clean this up, we choose a basis where
$\oint_{A^{k}}\alpha_{+}^{i}$ $=\delta^{i}_{k}$
$\oint_{B_{k}}\alpha_{+}^{i}$ $={\cal T}^{ik},$ (200)
and
$\oint_{A^{k}}\alpha_{-}^{i}$ $=\delta^{i}_{k}$
$\oint_{B_{k}}\alpha_{-}^{i}$ $=\tilde{\cal T}^{ik},$ (201)
so that, using a vector notation,
$Z_{z.m.}={\rm exp}\Bigg(-\frac{i}{2\pi}\sum_{\hat{\bf n},\hat{\bf m}}\hat{\bf m}\cdot(\tilde{\cal T}^{T}-{\cal T})\hat{\bf n}-$ $\frac{i}{2\pi}\sum_{\hat{\bf m}}\hat{\bf m}\cdot\left({\cal T}\oint_{\bf A}{\cal J}-\oint_{\bf B}{\cal J}\right)$
$+\frac{i}{2\pi}\sum_{\hat{\bf n}}\hat{\bf n}\cdot\left(\tilde{\cal T}\oint_{\bf A}{\cal J}-\oint_{\bf B}{\cal J}\right)\Bigg).$ (202)
However, we would like to normalise the periods to
$\oint_{A^{k}}dP=2\pi Rm_{k}\qquad\oint_{B_{k}}dP=2\pi Rn^{k},$ (203)
with $m_{k},n^{k}\in{\mathbb{Z}}$ and $R$ a fixed constant. With our normalization we find
$\hat{\bf m}=-2\pi R({\bf m}\tilde{\cal T}-{\bf n})({\cal T}-\tilde{\cal T})^{-1}\qquad\hat{\bf n}=2\pi R({\bf m}{\cal T}-{\bf n})({\cal T}-\tilde{\cal T})^{-1}\$(204)
so that $Z_{z.m.}$ becomes
$Z_{z.m.}$ $={\rm exp}\Bigg(-{2\pi iR^{2}}\sum_{{\bf m},{\bf n}}({\bf m}\tilde{\cal T}-{\bf n})({\cal T}-\tilde{\cal T})^{-1}\cdot({\cal T}-\tilde{\cal T}^{T})({\bf m}{\cal T}-{\bf n})({\cal T}-\tilde{\cal T})^{-1}$
$\qquad\qquad+{2iR}({\bf m}\tilde{\cal T}-{\bf n})({\cal T}-\tilde{\cal T})^{-1}\cdot\Big(\frac{1}{2}({\cal T}+\tilde{\cal T})\oint_{{\bf A}}{\cal J}-\oint_{{\bf B}}{\cal J}\Big)\Bigg).$ (205)
7 Discussion and Conclusions
In this paper we have studied the bi-metric generalisation Hull:2023dgp of the Sen formulation Sen:2015nph ; Sen:2019qit applied to chiral bosons on a Riemann surface. Analytically continuing these two metrics to Riemannian ones, we showed that the bi-metric action reduces to that of a non-chiral boson coupled to a complex metric. We were then able to compute the partition function and gave an explicit expression for the case in which the 2d-space is a torus.
Let us review our calculation. Formally, the Sen action is defined in terms of two Lorentzian metrics $\bar{g}$ and $g$ which we have parameterised by four independent real functions $\rho,\tilde{\rho},\tau,\tilde{\tau}$ with $\rho\neq\tilde{\rho}$ and $\tau\neq\tilde{\tau}$ (as well as conformal factors that drop out of our calculations). As such, the action is real and the partition function, as defined by the path-integral in (3.1), is purely oscillatory and ill-defined. However, we can analytically continue $\rho,\tilde{\rho},\tau,\tilde{\tau}$ to the complex plane where the action becomes complex and, at least for a certain range of $\rho,\tilde{\rho},\tau,\tilde{\tau}$, it is damped allowing us to evaluate the partition function. In particular, this includes the regime $\tilde{\rho}=\rho^{*}$ and $\tilde{\tau}=\tau^{*}$, where $\bar{g}$ and $g$ are Riemannian metrics, provided we take $\tau_{2}>0$ and $\rho_{2}<0$ and sufficiently large in magnitude.
We also discussed modular properties of our partition function and showed that in general it is only invariant under transformations where the two complex structures transform in the same way. This is expected since only these transformations arise from coordinate transformations acting on the torus. In particular the zero mode contributions to the partition function do not factorise. However, for applications to the heterotic string theory we are interested in partition functions where $\tau$ and $\rho$ can be independently transformed. We showed that by considering $n$ chiral scalars, with $n$ a multiple of 8, the corresponding partition function arising from a Narain Lattice can factorise into the simple form
$Z(\tau,\rho)=F(\tau)\tilde{F}(\rho)\ ,$ (206)
and hence one can construct theories where the modular transformations of $\tau$ and $\rho$ can be made independent and moreover such that the theory is invariant under two sets of modular transformations. We showed how to construct a consistent worldsheet heterotic string theory by including anomaly cancelling matter fields in both the physical and shadow sectors.
We have also seen that in order to have a well-defined partition function we needed to have $M\neq 0$ to provide the damping in the integrals. Specifically, for a given physical metric $g$, we chose the metric $\bar{g}$ so that $M\neq 0$. If we were to try to directly consider the case $M=0$, then the action would be simply
$S=\int Q^{\prime}\wedge dP,$ (207)
where $Q^{\prime}=2Q+\tfrac{1}{2}P+\tfrac{1}{2}\bar{\star}dP$. In two dimensions, this is a well-known conformal field theory; the holomorphic $\beta\gamma$ system, with $c=2$. However, it generalises to higher dimensions where it resembles a topological $B\wedge F$ theory except that there is now a metric dependent constraint $Q^{\prime}=\bar{\star}Q^{\prime}$. We have examined this theory separately in Hull:2025bqo .
Acknowledgements
We would like to thank Ashoke Sen for helpful discussions. The research of CH was supported by the STFC Consolidated Grant ST/X000575/1. N.L. is supported by the STFC grant ST/X000753/1.
Appendix A Appendix: Conventions
In this paper we use two independent metrics $\bar{g}$ and $g$ and their associated Hodge-star operators are $\bar{\star}$ and $\star$. Furthermore, to avoid any confusion we denote complex conjugation by .
We adopt the following convention: the zwiebein indices are labelled as $a,b=+,$-. In particular we never use ++ or - for world indices. For example
$V_{\mu}$ $=\bar{e}_{\mu}{}^{a}V_{a}$ (208)
$=\bar{e}_{\mu}{}^{+}V_{+}+\bar{e}_{\mu}{}^{-}V_{-},$ (209)
and
$\partial_{+}=\bar{e}_{+}{}^{\mu}\partial_{\mu}\qquad\partial_{-}=\bar{e}_{-}{}^{\mu}\partial_{\mu},$ (210)
which in general are not derivatives with respect to a pair of coordinates $x^{+}$ and $x^{-}$.
Next we choose a convention where $\epsilon_{\mu\nu}$ is defined to be anti-symmetric with $\epsilon_{01}=1$. Similarly $\epsilon^{\mu\nu}$ is anti-symmetric with $\epsilon^{01}=-1$. This implies that $\epsilon_{+-}=-\epsilon^{+-}=1$.
Given a metric, say $\bar{g}$, and one-form $V=V_{\mu}dx^{\mu}$ we define
$\bar{\star}V=\sqrt{-\bar{g}}\epsilon_{\mu\nu}\bar{g}^{\nu\lambda}V_{\lambda}dx^{\mu}.$ (211)
It follows that $\bar{\star}^{2}V=V$ for any one-form $V$ and for any metric $\bar{g}$, including complex ones.
The benefit of the zwiebein basis is that
$\bar{\star}Q_{a}=\epsilon_{ab}\eta^{bc}Q_{c},$ (212)
i.e.
$\bar{\star}Q_{\pm}=\pm Q_{\pm},$ (213)
and hence a condition such as $Q=\bar{\star}Q$ simply implies that $Q_{-}=0$.
Lastly for a two-form $\omega^{(2)}=\tfrac{1}{2}\omega_{\mu\nu}dx^{\mu}\wedge dx^{\nu}$ we define
$\int\omega^{(2)}=\frac{1}{2}\int\omega^{(2)}dx^{\mu}\wedge dx^{\nu}=-\frac{1}{2}\int\omega^{(2)}_{\mu\nu}\epsilon^{\mu\nu}d^{2}x.$ (214)